### v1

# FRAbench and GenEval: Scaling Fine-Grained Aspect Evaluation across Tasks, Modalities

Abstract
Evaluating the open-ended outputs of large language models (LLMs) has become a bottleneck as model capabilities, task diversity, and modality coverage rapidly expand. Existing “LLM-as-a-Judge” evaluators are typically narrow in a few tasks, aspects, or modalities, and easily suffer from low consistency. In this paper, we argue that explicit, fine-grained aspect specification is the key to both generalizability and objectivity in automated evaluation. To do so, we introduce a hierarchical aspect taxonomy spanning 112 aspects that unifies evaluation across four representative settings — Natural Language Generation, Image Understanding, Image Generation, and Interleaved Text-and-Image Generation. Building on this taxonomy, we create FRAbench, a benchmark comprising 60.4k pairwise samples with 325k aspect-level labels obtained from a combination of human and LLM annotations. FRAbench provides the first large-scale, multi-modal resource for training and meta-evaluating fine-grained LMM judges. Leveraging FRAbench, we develop GenEval, a fine-grained evaluator generalizable across tasks and modalities. Experiments show that GenEval (i) attains high agreement with GPT-4o and expert annotators, (ii) transfers robustly to unseen tasks and modalities, and (iii) reveals systematic weaknesses of current LMMs on evaluation [^1] Our code can be found at https://github.com/gsfsdv/FRABench-and-GenEval.
1 Introduction
As large language models (LLMs) have shown amazing abilities in human-like question answering and creative generation [1, 2], assessing the quality of their free-form outputs has become increasingly challenging. Automated evaluators, a.k.a. “LLM-as-a-Judge” paradigm [3], have therefore received much research attention, with the purpose of providing absolute or relative scores sometimes accompanied by detailed feedback [4]. While existing evaluators [5, 6] are typically tailored to limited tasks, aspects, or modalities, which sharply constrains their scope of application. Equally troubling, the credibility of these evaluators remains in doubt — empirical evidence shows that even human annotators suffer from a low agreement in evaluating the open-ended responses of LLMs as task complexity rises, and current automated evaluators inherit the same vulnerability.
In this paper, we target a fine-grained evaluator designed to generalize across different tasks and modalities (i.e., text and images). We highlight that aspect-level assessment is the core of such generalizability. On one hand, aspects like engagement, naturalness, or creativity are inherently interrelated [7], thus transferring similar semantics and evaluation standards for various tasks. On the other hand, explicit aspect specification can reduce subjectivity. Our observations indicate that the low agreement among human annotators is mainly caused by their implicitly varying weighting over aspects, rather than disagreement over the definition of any single aspect. For example, when judging story generation, some prioritize creativity while others emphasize logical consistency. By specifying aspects, one can substantially mitigate this variance source and obtain more objective evaluations.
Method Task Modalty Aspect Generalizable Data Source
NLG IU IG ITIG Text Image
X-Eval [5] 27 Human
Themis [8] 50 Synthetic
ImageReward [9] 3 Human
VisionReward [10] 37 Human
LLaVA-Critic [11] Synthetic
GenEval (ours) 112 Synthetic & Human
Table 1: Comparison of our GenEval with recent evaluators. GenEval supports evaluating across four tasks and two modalities, incorporates multiple aspects, and demonstrates strong generalization ability. “–” indicates that the number of supported aspects is not explicitly specified.
To this end, we first build a fine-grained aspect tree by conducting a comprehensive survey on four types of (multi-modal) tasks: Natural Language Generation (NLG), Image Understanding (IU), Image Generation (IG), and Interleaved Text-and-Image Generation (ITIG). We then manually organize, extend, and redefine 112 aspects with hierarchical relations as a universal evaluation taxonomy—aspect tree. Note that although recent studies like Prometheus [12] and BIGGen-Bench [13] use auto-generated aspects for ultra-fine evaluation, we avoid this approach for two reasons. First, when the aspect becomes that narrowly tailored, its boundaries with direct evaluation become blurred, thereby reintroducing the very subjectivity we aim to suppress. Second, the universe of evaluation aspects is comparatively stable and, in practice, enumerable. Once the application scenario is specified, human or automated evaluators can simply traverse our aspect tree to select the relevant ones. This design dramatically lowers the cognitive load of defining evaluation standards and promotes consistency across tasks and modalities.
Based on the aspect tree, we construct a large-scale dataset named Fine-grained Aspect Benchmark (FRABench) for both training and testing. We leverage the aspect tree to select relevant multiple aspects for each of the 60.4k pairwise responses across 28 sub-tasks. To obtain high-quality evaluation labels while maintaining cost efficiency, we collect available human annotations wherever possible and supplement them with GPT-4o-assisted completions. In total, this process yields 325k evaluation labels. Using FRABench, we develop GenEval, a generalizable evaluator. Table 1 shows the detailed comparison with existing evaluators. Experimental results demonstrate that GenEval exhibits strong generalization ability on both in-domain and out-of-domain tests, and achieves high consistency with GPT-4o and human annotations. Meanwhile, we also identify several limitations and deficiencies in current LMMs when evaluating IG tasks, which can serve as promising directions for future research.
In summary, our contributions are as follows: (1) We construct an evaluation aspect tree covering 112 aspects to ease fine-grained evaluation. (2) We develop FRABench, the first large-scale, multi-modal resource for training and meta-evaluating fine-grained LLM judges. (3) We create GenEval and conduct experiments to showcase its generalization capabilities, providing key insights.
2 Fine-grained Aspect Tree Construction
To ease fine-grained evaluation for both automated evaluators and humans, we construct an aspect tree that is universal across modalities and tasks. The construction pipeline involves two main steps: (1) Collecting and extending aspects. (2) Constructing aspect taxonomy.
2.1 Aspect Collection and Extension
To fully leverage existing aspects, we first collect 28 sub-tasks under the four types of tasks (i.e., NLG, IU, IG, and ITIG). These sub-tasks span all six combinations of input types (text and text-with-image) and output types (text, image, and text-with-image), ensuring comprehensive coverage across multimodal tasks. Subsequently, we systematically collected literature related to each sub-task to gather relevant evaluation aspects and, where available, their definitions. For aspects without available definitions, we define them according to their practical meaning.
[…]
2.2 Aspect Taxonomy Construction
Figure 1: Universal Aspects: A subtree rooted at overall, driven by task independence.
Figure 2: Task-specific Aspects: A subtree rooted at overall, driven by task dependence.
In this section, we construct an aspect tree based on the aspects to establish a standardized evaluation taxonomy for multimodal evaluation tasks. Based on this universal taxonomy, both human and automated evaluators can easily traverse the aspect tree to customize evaluation criteria for specific tasks or scenarios, enabling fine-grained and objective assessments.
We first use the overall aspect as the root node of the aspect tree, and then divide the remaining aspects into two subtrees based on their task independence and task specificity:
  • Universal Aspects (UAs): These aspects exhibit task independence, as they primarily focus on assessing the quality of the model’s output. They are often modality-specific (for example, fluency for text or fidelity for image) and serve as fundamental quality aspects across tasks.
  • Task-specific Aspects (TAs): These aspects exhibit task dependence, as they primarily focus on task completion rather than output quality. They are typically closely tied to the task type. For example, engagingness for story generation or accuracy for mathematical reasoning.
For the hierarchical construction within the two subtrees, we first obtain existing tree structures from previous studies [14, 15, 16, 17, 18, 19]. We then classify these hierarchical structures, together with aspects lacking predefined hierarchies, into UAs and TAs as described above. Finally, for aspects without predefined hierarchies, we determine their positions within the tree according to their definitions—either assigning them to an appropriate location within an existing tree or establishing them as new root nodes. For example, the definition of alignment encompasses both object alignment and attribute alignment. This definition-driven approach is reasonable, as it ensures that the constructed hierarchy accurately reflects the semantic relationships among aspects. The final two subtrees are shown in Figure 2 and Figure 2.
Based on the aspect tree, the aspect-level evaluation process can be streamlined. Specifically, we first determine the correspondence between each UAs subtree (e.g, the subtree rooted at readability, represented by readability) and its respective modality: readability for text, image quality for image, image coherence for multi-images, and text-image relationship for text-with-image, while bias and toxicity apply across all modalities. Once the evaluation task is determined, we can identify the corresponding subtree based on the output modality to select appropriate aspects. Additionally, task-specific aspects can be traversed from the TAs according to the objectives of the task.
3 FRABench Construction
To support fine-grained evaluation, we construct FRABench, a comprehensive dataset for training and evaluating evaluators on the aspect tree we proposed. Following prior studies [20, 21] that demonstrate pointwise scoring is more susceptible to contextual bias, we adopt a pairwise comparison evaluation method in FRABench. FRABench comprises 60.4K pairwise samples and 325k evaluation labels. Detailed statistics are presented in Figure 3 and Appendix B.4. Moreover, we rigorously partition the FRABench dataset into training, in-domain test, and out-of-domain test subsets to enable evaluation of both in-domain and out-of-domain generalization capabilities of evaluators.
Source Tasks Dataset Sub-Tasks Split Size
Public NLG $\blacksquare$ Auto-J Summarization, Creative Writing, Rewriting, etc. Train / FRA-ID 2.5k / 0.6k
Question Generation, Title Generation, etc. FRA-OOD 0.6k
IU $\blacksquare$ VLFeedback Image Captioning, Text-rich Understanding, etc. Train / FRA-ID 14.5k / 2.4k
Medical VQA, Academic VQA, etc. FRA-OOD 1.2k
IG $\blacksquare$ GenAI-Bench Text-to-Image Generation FRA-OOD 1.4k
$\blacksquare$ ImageRewardDB Text-to-Image Generation Train / FRA-ID 6k / 1.5k
Generated IU $\blacksquare$ TextVQA Text Reasoning Train / FRA-ID 2.1k / 0.4 k
$\blacksquare$ ChartVQA Chart Reasoning Train / FRA-ID 2.1k / 0.4 k
$\blacksquare$ InfographicsVQA Graph Reasoning Train / FRA-ID 2.1k / 0.4 k
IG $\blacksquare$ MagicBrush Image Editing Train / FRA-ID 3.5 k / 0.5 k
$\blacksquare$ COCO Text-to-Image Generation Train / FRA-ID 4.8k / 1k
ITIG $\blacksquare$ VIST Visual Story Completion Train / FRA-ID 5.5k / 0.7k
$\blacksquare$ wikiHow Multimodal Script Generation Train / FRA-ID 6.4k / 0.6k
$\blacksquare$ InterleavedBench Storytelling Generation, Activitynet Continuation FRA-OOD 90
Figure 3: Data statistics of the FRAbench. Besides using public datasets, we generated an equivalent amount of pairwise data. Detailed counts of evaluation labels for each sub-task are in Appendix B.4.
3.1 Dataset Collection
We collect queries from the sub-tasks in the aspect tree we proposed and generate pairwise samples. Specifically, we obtain 29.3K response pairs from high-quality public datasets and generate the remaining 30.1K pairs using LMMs (model details provided in Appendix B.1). After obtaining the queries and response pairs, we assign aspects from our aspects tree(the aspect selection method is described in Section 2.2). As a result, each pairwise sample is annotated with an average of 8 UAs and 3 TAs. We then generate fine-grained evaluation labels for each response pair. Our evaluation labels have two sources: (1) Human annotations: We directly obtain human-annotated evaluation scores covering three IG-related aspects from the ImageRewardDB dataset [9], and supplement corresponding feedback using GPT-4o to form the evaluation labels. (2) GPT-4o annotations: For all other pairwise samples, due to the absence of human-annotated labels, we use GPT-4o to generate evaluation labels. All template details are provided in Appendix F. Moreover, to mitigate position bias commonly observed in evaluators [21], we balance the number of samples where response 1 is preferred over response 2 and vice versa. Finally, we generate 325k fine-grained evaluation labels.
3.2 Dataset Partitioning
To comprehensively train and evaluate the generality of GenEval, we split FRABench into training, in-domain, and out-of-domain test subsets. Specifically, we randomly select 18 sub-tasks from 4 tasks, covering 22 UAs and 35 TAs for the training set and in-domain test set (FRA-ID). Meanwhile, we select 10 unseen sub-tasks covering 28 seen UAs and 27 unseen TAs for the out-of-domain test set (FRA-OOD). Additionally, to assess GenEval’s generalization across UAs, we construct FRAUAs-OOD, which includes the same sub-tasks as the training set but different UAs across 4 tasks. Finally, the training set contains 255.4K samples, FRA-ID contains 45.2K samples, FRA-OOD contains 24.4K samples, and FRAUAs-OOD contains 5.3K samples.
The role of an evaluator is to assess model responses in place of humans. Therefore, the primary criterion for evaluating an evaluator’s effectiveness is their alignment with human judgments. Therefore, we construct corresponding evaluation datasets to measure the consistency between the evaluator’s assessments and human judgments. We extract partial samples from FRA-ID and FRA-OOD, and recruited three human annotators with master’s degrees or higher for annotation. Finally, we retain 6.9K in-domain samples and 6.0K out-of-domain samples with highly consistent human annotations, forming the final in-domain evaluation dataset (FRA-ID-H) and out-of-domain evaluation dataset (FRA-OOD-H), respectively. Details of the human annotation process are provided in Appendix E.
4 Evaluator Training
To train the GenEval, we use the training set from FRABench with Qwen2-VL-7B-Instruct as the backbone. As our primary focus is on fine-grained evaluation and the construction of the aspect tree and FRABench, we adopt the SFT for training. More advanced training techniques, such as RL, are left for future work. For training strategies, we explore the effectiveness of staged training versus unified training in the multi-task setting. Specifically, we split the training data by output modality (text, image, and text-image) and perform staged training in all six possible orders, producing six evaluators. The results indicate that phased training causes catastrophic forgetting. As a result, we adopt the unified training strategy. Details of the training setup and the results of the phased training experiments are presented in Appendix C.1 and Appendix D.2, respectively.
5 Experiments
To comprehensively evaluate GenEval, we consider three experimental settings: In-Domain Evaluation, Out-of-Domain Evaluation, and Evaluation as an Evaluator and Reward Model. Next, we present the experimental setup and corresponding results.
5.1 Experimental Setup
Benchmarks. (1) In-Domain Evaluation: we use FRA-ID and FRA-ID-H for assessment. (2) Out-of-Domain Evaluation: FRAUAs-ODD, FRA-OOD, and FRA-OOD-H are employed. We also include some public benchmarks: SummEval [22] and MANS [23] for NLG, WildVision [24] for IU, and GenAI-Bench [25] and Winoground [26] for IG. (3) Evaluation as Evaluator and Reward Model: MT-Bench [27] and MLLM-as-a-Judge [28] are used to evaluate models as an evaluator. VLRewardBench [29] is used for reward model evaluation. Detailed descriptions of each benchmark are provided in Appendix C.2.
Baselines. we use following models as baselines: GPT-4o, Claude-3.5, Qwen2-VL-72B-Instruct (Qwen2VL-72B), Qwen2-VL-7B-Instruct (Qwen2VL-7B), Themis-8B [8] (Themis), LLaVA-Critic-7B [11] (LLaVA-Critic), ImageReward [9] and VisionReward [9]. Notably, the paradigm adopted by ImageReward and VisionReward for fine-grained evaluation differs from ours. Therefore, they are used solely as baselines in public benchmark evaluations.
5.2 In-Domain Evaluation
For in-domain evaluation, we assess performance using the FRA-ID and FRA-ID-H. Figure 4 presents radar charts that illustrate the overall degree of alignment [^3] We use aspect abbreviations here; the full names can be found in Appendix A.. Detailed scores for each aspect are reported in Appendix D.1. The accompanying Table 2 reports average accuracies for each task. Additionally, we also sample three types of aspects for aspect-level analysis. As shown in Figure 5, we select aspects from four test sets based on GenEval’s performance: (1) Weak aspects, where the best model surpasses GenEval by over 20%. (2) Universal poor aspects, where scores from all evaluators are very low. (3) Strong aspects, where GenEval excels. Only two weak aspects appear across all test sets, and both are included. For the other categories, we choose representative subsets.
The results demonstrate that GenEval consistently achieves higher alignment with both GPT-4o and human annotators compared to the baselines (see shaded area in radar chart). Specifically, excluding GPT-4o, GenEval surpasses all baselines on NLG, IU, and ITIG tasks within FRA-ID and FRA-ID-H, achieving overall average accuracies of 88.7% and 74.6%, respectively. On NLG, GenEval outperforms Themis, while our base model Qwen performs much worse, suggesting that with appropriate training, LMM-based evaluators can also effectively assess NLG tasks. Moreover, GenEval shows strong performance on the ITIG of FRA-ID-H, where its alignment with human annotators exceeds that of GPT-4o. This improvement can be attributed to GenEval’s generalization ability and cross-task transfer performance. For aspect-level comparison, GenEval outperforms GPT-4o on some aspects, with two illustrative examples shown in Figure 5a.
Figure 4: Comparison on FRA-ID and FRA-ID-H. The top shows alignment with GPT-4o using FRA-ID, the bottom shows alignment with human annotators using FRA-ID-H. Each point is the average accuracy (with ties) for an aspect shared across sub-tasks of the same task.
Method Seen UAs for Seen Tasks Seen TAs for Seen Tasks
GPT-4o (FRA-ID) Human (FRA-ID-H) GPT-4o (FRA-ID) Human (FRA-ID-H)
NLG IU IG ITIG NLG IU IG ITIG NLG IU IG ITIG NLG IU IG ITIG
GPT-4o 94.0 91.7 85.5 90.9 76.1 79.1 65.0 66.1 92.5 90.0 76.6 89.4 79.5 74.8 58.8 73.6
Claude-3.5 79.4 79.1 55.2 77.2 67.2 71.9 52.5 61.1 80.3 78.0 52.7 86.4 76.3 64.8 57.8 75.9
Qwen2VL-72B 78.1 74.0 70.0 58.5 70.1 70.5 67.5 56.9 76.1 68.8 53.4 75.4 79.1 67.4 47.8 70.3
Qwen2VL-7B 51.1 50.0 58.4 37.4 48.6 55.6 57.7 38.2 48.9 46.9 42.6 39.5 46.7 51.6 34.6 36.1
Themis 60.4 - - - 60.3 - - - 47.5 - - - 53.4 - - -
LLaVA-Critic - 47.5 - - - 39.0 - - - 68.0 - - - 56.0 - -
Ours 91.1 88.2 61.0 87.0 75.8 78.2 65.0 70.0 87.9 85.1 71.9 87.2 84.1 71.5 58.1 73.9
Table 2: Average accuracy on FRA-ID and FRA-ID-H, separated by UAs and TAs. Average accuracy (with ties) is calculated by averaging the evaluation accuracies across sub-tasks within each task.
Figure 5: Comparison of aspect-level. In the subfigures, the x-axis shows the test set (FRA-ID, FRA-ID-H, FRA-OOD, or FRA-OOD-H) and task (NLG, IU, IG, or ITIG) of each evaluated aspect.
We also observe special phenomena. As shown in Table 2, IG has the lowest accuracy among the four tasks across all baselines. In aspect analysis, action alignment and object alignment in IG under FRA-ID-H receive consistently low scores across all baselines (Figure 5b). We will discuss the reasons in Section 5.3.1. Additionally, for the harmfulness of IG in FRA-IN (Figure 5c), GenEval demonstrates relatively poor performance. Our analysis indicates that this issue mainly arises because GenEval classifies shadowy or gloomy images as harmful due to its heightened sensitivity to elements that could evoke psychological discomfort. Relevant cases are presented in Appendix G.
Figure 6: Comparison on FRA-OOD and FRA-OOD-H. Black sectors show unseen TAs (for task and aspect generalization), while the rest show seen UAs for unseen tasks (for task generalization).
Method Seen UAs for Unseen Tasks Unseen TAs for Unseen Tasks
GPT4o (FRA-OOD) Human (FRA-OOD-H) GPT4o (FRA-OOD) Human (FRA-OOD-H)
NLG IU IG ITIG NLG IU IG ITIG NLG IU IG ITIG NLG IU IG ITIG
GPT-4o 90.8 94.7 89.9 88.2 84.0 82.1 72.3 93.1 87.3 89.7 85.9 91.0 83.2 82.1 74.2 93.1
Claude-3.5 74.6 85.8 72.5 75.6 83.0 76.5 63.1 91.0 84.1 84.3 65.6 85.0 82.6 76.5 65.1 91.0
Qwen2VL-72B 70.2 82.4 65.8 60.0 78.3 75.3 48.6 83.7 76.3 81.5 65.6 85.0 77.3 75.3 53.8 83.7
Qwen2VL-7B 50.4 65.9 61.4 43.4 50.9 65.9 40.9 44.3 54.5 69.1 37.7 69.1 49.1 65.9 46.0 44.3
Themis 56.7 - - - 58.9 - - - 55.0 - - - 58.8 - - -
LLaVA-Critic - 52.2 - - - 76.2 - - - 80.8 - - - 76.2 - -
Ours 81.7 90.4 69.0 83.1 79.0 80.9 62.1 90.6 83.0 86.3 62.9 89.3 78.3 80.9 66.1 90.6
Table 3: Average accuracy on FRA-OOD and FRA-OOD-H. Seen UAs for Unseen Tasks are used to assess task generalization, while Unseen TAs for Unseen Tasks evaluate aspect generalization.
5.3 Out-of-Domain Evaluation
5.3.1 Tasks Generalization Evaluation
We use FRA-OOD and FRA-OOD-H to evaluate GenEval’s generalization to tasks. The results are shown in the regions outside the black sectors in Figure 6. GenEval demonstrates strong alignment with both GPT-4o and human annotators on all tasks. In the “Seen UAs for Unseen Tasks” column of Table 3, GenEval achieves overall average accuracies of 85% and 83% on FRA-OOD and FRA-OOD-H, respectively, and outperforms both Themis and LLaVA-Critic on all tasks. Supporting the effectiveness of aspect-level evaluation for cross-task generalization. As illustrated in Figure 5a for aspect-level comparison, GenEval surpasses the baselines across most aspects, with only five aspects displayed here. We also provide several good cases in the Appendix G.
However, as shown in Table 3, results for IG are consistent across in-domain evaluation, with all baselines performing poorly. As shown in Figure 5b, all evaluators perform poorly on fidelity. Our analysis indicates that these aspects in IG require recognizing fine-grained object features, such as well-formed facial features in humans, which models fail to detect. GPT-4o exhibits similar limitations (detailed samples are shown in Appendix G). This suggests that, although the evaluator learns a broad range of aspects, it still lacks a better understanding of objective visual semantics.
5.3.2 Aspects Generalization Evaluation
We use the unseen TAs in FRA-OOD and FRA-OOD-H to evaluate GenEval’s generalization to aspects. To support the validity of aspect generalization evaluation, we provide ROUGE_L [30] tests on unseen aspects in Appendix B.2. The results are shown in the black sectors of Figure 6, and in Table 3. In terms of overall coverage, GenEval outperforms both Themis and LLaVA-Critic, the state-of-the-art evaluators in their respective domains. It continues to demonstrate strong overall performance on the NLG, IU, and ITIG, with high consistency with both GPT-4o and human annotations, achieving average accuracies of 86.2% and 83.2%, respectively. For aspect-level comparison, as shown in Figure 5a, GenEval outperforms GPT-4o on unseen aspects such as suitability and explainability under NLG and IU in FRA-OOD-H. Additionally, for UAs generalization evaluation, since GenEval has already transferred relevant UAs from other tasks, it performs well on FRAUAs-OOD. Experimental results are provided in Appendix D.1. These results demonstrate that even without exposure to these unseen aspects during training, GenEval remains effective in assessment. This is largely due to the inherently interrelated nature of aspects, which enables the transfer of similar semantics and evaluation standards across different tasks.
5.3.3 Public Benchmark Evaluation
Additionally, to verify the reliability and generalization of GenEval, we evaluate it on some public meta-evaluation benchmarks. Tables 4, 6, and6 show results for the NLG, IU, and IG tasks, respectively. GenEval outperforms Themis (NLG) and LLaVA-Critic (IU), both task-specific evaluators, indicating that our cross-task evaluator with aspect-level evaluation performs better than single-task evaluators. For IG tasks, GenEval performs on par with VisionReward and surpasses ImageReward. Overall, the results show that training on cross-modality and cross-task can lead to stronger performance than training on a single task or modality.
Method SummEval MANS
Coherence Consistency Fluency Relevance Ave. diff($\uparrow$)
tau($\uparrow$) diff($\uparrow$) tau($\uparrow$) diff($\uparrow$) tau($\uparrow$) diff($\uparrow$) tau($\uparrow$) diff($\uparrow$) tau($\uparrow$) diff($\uparrow$)
GPT-4o 58.0 64.2 79.1 85.1 64.3 72.8 60.1 67.1 65.3 72.3 68.5
Claude-3.5 63.5 70.6 81.6 87.9 73.7 83.1 59.6 66.6 69.6 77.0 68.4
Qwen2VL-72B 66.8 73.2 66.8 66.4 71.8 80.4 62.2 69.3 66.9 72.3 19.3
Qwen2VL-7B 48.8 52.5 35.5 30.6 44.8 49.8 41.7 44.5 42.7 44.3 60.2
Themis 60.7 62.1 81.8 86.0 73.3 77.7 54.4 54.6 67.5 70.1 44.2
Ours 64.6 71.8 75.2 83.5 74.7 84.5 61.3 67.6 69.0 76.9 69.3
Table 4: The results of GenEval compared with baselines on two benchmarks. “tau” indicates that accuracy is calculated with ties, and “diff” excludes tied pairs when calculating accuracy.
Method WildVision
tau ($\uparrow$) diff ($\uparrow$) $\tau$ ($\uparrow$)
GPT-4o 55.3 70.1 73.3
Claude-3.5 53.3 67.3 61.2
Qwen2VL-72B 67.3 59.6 65.5
Qwen2VL-7B 39.2 40.6 23.1
LLaVA-Critic 53.0 66.0 59.6
Ours 53.9 68.6 66.5
Table 5: For comparison in wild settings. $\tau$ indicates the comparison of model rankings using Elo ratings.
Method GenAI-Bench Winoground
tau($\uparrow$) diff($\uparrow$) Relation Object Both Ave.
diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$)
GPT-4o 55.6 69.5 62.6 73.0 73.0 69.5
Claude-3.5 55.6 71.0 71.2 73.0 69.2 71.1
Qwen2VL-72B 49.1 52.6 46.7 60.9 57.6 55.0
Qwen2VL-7B 35.8 38.0 33.4 42.5 46.1 40.6
VisionReward 51.0 66.4 60.2 64.1 74.9 66.4
ImageReward 48.6 64.9 54.0 58.2 69.2 60.4
Ours 53.6 65.5 57.5 59.1 80.7 65.7
Table 6: Comparison of GenEval and other baselines on two text-to-image meta-evaluation benchmarks.
5.4 Evaluation as Evaluator and Reward Model
In this section, we assess GenEval’s performance as an evaluator on the MT-Bench and MLLM-as-a-Judge benchmarks. We also assess its potential as a multimodal reward model using VLRewardBench. As shown in Table 8 and Table 8, GenEval outperforms Themis—the leading evaluator in the LLM-as-a-Judge setting. For MLLM-as-a-Judge, GenEval also demonstrates modest advantages over LLaVA-Critic and surpasses it in VLRewardBench. These results indicate that GenEval serves effectively as an evaluator for both NLG and IU tasks and further verify the potential of FRABench as training data for multimodal reward models.
Method MT-Bench
tau($\uparrow$) diff($\uparrow$)
GPT-4o 70.9 83.5
Claude-3.5 76.3 90.7
Qwen2VL-72B 75.9 88.7
Qwen2VL-7B 44.5 50.1
Themis 43.6 37.7
Ours 74.9 88.3
Table 7: Test results of the baselines as evaluators.
Method MLLM-as-a-Judge VLRewardBench
tau ($\uparrow$) diff ($\uparrow$) General Hallucination Reasoning Ave.
diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$)
GPT-4o 58.1 67.0 50.2 81.4 74.8 68.8
Claude-3.5 58.4 68.3 38.5 82.6 66.1 62.4
Qwen2VL-72B 54.6 58.5 50.8 75.4 70.7 65.6
Qwen2VL-7B 41.3 44.6 45.1 62.8 62.5 56.8
LLaVA-Critic 55.6 65.5 42.0 41.2 60.0 47.7
Ours 57.2 67.0 46.4 57.7 71.1 58.4
Table 8: Test results of the baselines as multimodal evaluators and reward models on benchmarks.
6 Related Work
6.1 Coarse-grained Overall-oriented Evaluation
Previously, coarse-grained evaluators, which produce overall judgments based on one or several aspects, have been widely explored across various evaluation tasks [31, 32, 33, 34, 35, 36]. For instance, in NLG, Wang et al. [37] first introduce PandaLM, a fine-tuned LLM designed to evaluate pairwise texts based on several aspect. Similarly, Li et al. [38] propose Auto-J, an evaluator capable of handling a broader range of tasks and aspects, supporting both pointwise and pairwise evaluation settings. In image generation evaluation, Xu et al. [9] develop ImageReward, an evaluator trained on a large-scale human annotation dataset. Using alignment and fidelity as reference dimensions, ImageReward provides an overall assessment. Despite showing good overall performance, coarse-grained evaluations lack the granularity needed to diagnose specific model deficiencies and often introduce aspect bias into the evaluation process.
6.2 Fine-grained Aspect-level Evaluation
To address the limitations of coarse-grained evaluation, recent studies have shifted toward fine-grained evaluation [33, 39, 6, 12, 40, 41, 42], where LMs are fine-tuned on multi-aspect datasets to produce aspect-specific judgments. For example, in NLG, Hu et al.[8] propose Themis, an LLM trained on the GPT-4 annotated NLG-Eval corpus. In image understanding, LLaVA-Critic[11] is the first fine-grained evaluator integrating diverse criteria, showing strong correlation with GPT-4o. For image generation, Xu et al.[10] introduce VisionReward, a VQA-based evaluator that assesses image quality across fine-grained aspects. While these methods improve over coarse-grained evaluation by targeting specific aspects, they struggle with scalability across different tasks and aspects. Liu et al.[5] further investigate aspect generalization through X-Eval, a two-stage learning framework that incorporates auxiliary aspects and demonstrates partial generalization in NLG. However, its reliance on predefined reference aspects and lack of open-source implementation restricts its usability and reproducibility.
7 Conclusion, Limitation and Future Work
This paper introduces GenEval, a generalizable fine-grained evaluator for cross-task and cross-modal evaluation. We start by building a comprehensive aspect tree that leads to the creation of FRABench, a large-scale evaluation dataset. GenEval is trained on diverse tasks and aspects to enhance its generalization capabilities. Experimental results demonstrate that GenEval achieves strong generalization through aspect-level evaluation. However, compared to the other three tasks, GenEval and other LMMs have relatively limited performance in image generation. This may primarily be due to existing LMMs’ insufficient active visual semantic understanding, which fails to autonomously identify vision defects. In future work, we plan to explore RAG-based methods to address LMMs’ lack of visual grounding and further evaluate GenEval’s effectiveness in generating DPO training data and its potential with reinforcement learning using FRABench.
Appendix A Aspect Sources and Definitions
We provide the sources and definitions of aspects under two categories: UAs and TAs, as shown in Tables 13 to 26. For UAs, the “Target” column indicates the applicable output modality: “T’ for Text, “I” for Image, and “ITI” for Interleaved Text-with-Image. For TAs, the “Sub-Task” column specifies the corresponding task type, as defined in Figure 3. Additionally, underscores are used to benchmark extended TAs for ITIG.
Task Sub-Task Split Size Task Sub-Task Split Size
NLG (Public) Summarization Train 218 IU (Public) Detailed Image Captioning Train 2.1k
Test 54 Test 0.4k
Creative Writing Train 575 Robustness-oriented Instructions Train 4k
Test 120 Test 0.4k
Rewriting Train 194 Medical Image Understanding Train 2.1k
Test 26 Test 0.4k
General Communication Train 1080 Text-rich Understanding Train 2.1k
Test 263 Test 0.4k
Functional Writing Train 433 General Visual Conversation Train 2.1k
Test 147 Test 0.4k
Question Generation Test 132 Simple Image Captioning Train 2.1k
Test 0.4k
Title Generation Test 38 Embodied Decision-making Test 0.4k
Keywords Extraction Test 80 Medical VQA Test 0.4k
Data Analysis Test 170 Academic VQA Test 0.4k
Translation Test 130 ITIG (Generated) Storytelling Generation Test 50
Activity Continuation Test 40
Table 9: The detailed statistical information of the sub-tasks is not shown in Figure 3
Task Sub-Task Model
IU Text Reasoning LLaVA-1.5-13B, Qwen2-VL-72B, InstructBLIP-Vicuna-13B, InternVL2-26B, Molmo-7B-D
Chart Reasoning
Graph Reasoning
IG Text-to-Image Generation Show-o-1.3B, Seed-X-17B, Flux-12B, Stable-Diffusion-3.5-Large, Ground Truth
Image Editin MagicBrush, SEED-X-17, InsPix2Pix, MGIE, Ground Truth
ITIG Visual Story Completion miniGPT, GILL, MM-Interleaved, GPT4o+SDXL, Ground Truth
Multimodal Script Generation
Storytelling Generation
Activitynet Continuation
Table 10: Different LMMs used for generating pairwise responses for sub-tasks under the “Generated” column in Figure 3
Appendix B Analysis of the FRABench
In this section, we provide detailed information and comprehensive analysis about the FRABench.
B.1 More Detailed Information about the Dataset
We provide the detailed information about sub-tasks for NLG and IU in Table 9, as this information is not shown in Figure 3. For the tasks in the “Generated” column in Figure 3, due to the absence of pairwise data, we use four LMMs with varying performance levels to generate pairwise responses for each query, as illustrated in Table 10.
B.2 Aspect Diversity Analysis
Following prior work [30, 43], we analyze the ROUGE-L distribution between aspects in the training set and the FRA-OOD, respectively. Specifically, we sample pairs of aspects from the training set and compute their ROUGE-L scores. The overall distribution is shown in Figure 7, which indicates that the selected aspects are distinct from one another, confirming the inclusion of diverse and novel aspects in the training set. Additionally, we measure the ROUGE-L scores between aspects by sampling one aspect from the training set and another from the FRA-OOD . As shown in Figure 8, aspects in the training set and those in the FRA-OOD do not overlap, demonstrating the validity of our out-of-domain evaluations.
Figure 7: Rouge-L score distribution among two randomly sampled aspects from the training set. A left-skewed distribution with low values shows that the aspects are diverse.
Figure 8: Rouge-L score distribution among a randomly sampled aspect from the training set and an aspect from the FRA-OOD. A left-skewed distribution with low values shows that they do not overlap with each other, hence meaning that the unseen aspect assumption is satisfied.
B.3 Position Bias Analysis
We provide the counts of evaluation samples in the training set, FRA-ID, FRAUAs-OOD, and FRA-OOD, showing cases where Response 1 is better than, equal to, or worse than Response 2, as illustrated in Figures 10 to 12. As shown in Figure 10, the four tasks in the training set maintain a relatively consistent distribution. In Table 11, we also evaluate the position consistency of GenEval on the FRA-ID and FRA-OOD. We perform two inferences on the same sample, swapping the positions of the responses in the second inference, and calculate the average consistency, i.e., whether the results from the two inferences align. These experimental results demonstrate that GenEval achieves higher positional consistency compared to other baselines.
Method FRA-ID FRA-OOD
NLG IU IG ITIG NLG IU IG ITIG
GPT-4o 81.4 85.2 78.3 81.9 76.3 83.0 75.7 86.0
Claude-3.5 79.9 75.1 77.4 85.5 74.9 72.1 75.4 88.5
Qwen2VL-72B 76.0 69.9 77.8 75.2 76.2 78.1 70.0 86.6
Qwen2VL-7B 29.7 27.4 62.4 19.6 24.5 30.9 46.5 20.5
LLaVA-Critic - 82.7 - - - 84.7 - -
Ours 82.1 82.6 75.8 83.8 78.9 88.2 75.5 89.8
Table 11: We calculate the models’ position consistency in the FRA-ID and FRA-OOD. Since our evaluator is trained on a relatively balanced dataset, it can mitigate the impact of position bias.
Figure 9: Statistics of pairwise data in the training set.
Figure 10: Statistics of pairwise data in the FRAUAs-OOD.
Figure 11: Statistics of pairwise data in the FRA-ID.
Figure 12: Statistics of pairwise data in the FRA-OOD.
B.4 Aspect Distribution Across Different Sub-tasks
We provide detailed information on the aspects selected for the training set, FRA-ID, FRAUAs-OOD, and FRA-OOD , as presented in Table 12. Additionally, we include statistics regarding the number of evaluation labels used for training and testing in each sub-task. Blue-marked aspects are human-annotated, while the rest are annotated by GPT-4o. The human-annotated data in IG is sourced from the ImageRewardDB dataset [9].
DS Sub-Task Aspect Split Universal Aspect Specific-task Aspect Size
Training Set & FRA-ID Summarization Train / Test Clarity, Grammaticality, Coherence, Fluency Coverage, Length Constraint, Layout, Conciseness 1.7k / 0.4k
Rewriting Train / Test Grammaticality, Coherence, Fluency, Clarity Accuracy, Pointing Out, Difference 1.3k / 0.1k
Creative Writing Train / Test Gender Bias, Ethnic Bias, Regional Bias Creativity, Engagingness, Theme 5.1k / 1k
Harmfulness, Grammaticality, Readability
Functional Writing Train / Test Clarity, Coherence, Fluency Relevance, Example Quality 3.4k / 1.1k
Gender Bias, Ethnic Bias Contextual Information
General Communication Train / Test Gender Bias, Ethnic Bias, Regional Bias Accuracy, Information Richness 11.8k / 2.6k
Clarity, Coherence, Fluency Understandability
General Visual Conv Train / Test Clarity, Harmfulness Helpfulness, Instruction Following 8.4k / 1.6k
Text-rich Und Train / Test Harmfulness, Vocabulary Complexity, Clarity Accuracy, Instruction Following 10.5k / 2k
Robustness-oriented Instructions Train / Test Spelling Accuracy, Structure Accuracy, Privacy Violations Instruction Inconsistency, Context Inconsistency 19.6k / 2.5k
Medical Image Und Train / Test Vocabulary Complexity, Bias, Sentence Complexity Accuracy 8.3k / 2k
Chart Reasoning Train / Test - Accuracy, Instruction Following 4.2k / 0.8k
Text Reasoning Train / Test - Accuracy, Instruction Following 4.2k / 0.8k
Graph Reasoning Train / Test - Accuracy, Instruction Following 4.2k / 0.8k
Simple Image Captioning Train / Test Readability, Bias Conciseness, Object Alignment 12.5k / 2.4k
Text-Image Relationship, Alignment
Detailed Image Captioning Train / Test Coherence, Fluency, Clarity, Bias Completeness, Attribute Alignment, Spatial Alignment 27.2k / 5.2k
Count Alignment, Action Alignment, Object Alignment
Color Alignment, Texture Alignment, Shape Alignment
Text-to-Image Generation (COCO&DB) Train / Test Harmfulness, Fidelity Count Alignment, Object Relationship 49.2k / 7.5k
Texture Alignment, Attribute Alignment,
Object Alignment, Spatial Alignment, Alignment
Action Alignment, Color Alignment, Shape Alignment
Image Editing Train / Test - Accuracy 3.5k / 0.5k
Visual Story Completion Train / Test Scene Consistency, Bias, Toxicity, Character Consistency Creativity, Engagingness, Completeness 39.8k / 7.7k
Stylistic Consistency, Action Consistency, Fluency
Multimodal Script Generation Train / Test Action Consistency, Coherence, Image Repetition Completeness, Feasibility 40.5k / 6.2k
Readability, Semantic Consistency, Stylistic Consistency Helpfulness, Safety
FRAUAs-OOD Summarization Test Sentence Complexity, Vocabulary Complexity, Readability - 1.3k
Spelling Accuracy, Structure Accuracy, Text Repetition
Rewriting Test Tense Consistency, Spelling Accuracy - 0.9k
Structure Accuracy, Text Repetition, Readability
Creative Writing Test Coherence, Fluency, Clarity, Text Repetition - 3.4k
Sentence Complexity, Vocabulary Complexity, Cultural Bias
Functional Writing Test Readability, Spelling Accuracy, Structure Accuracy - 2.1k
Sentence Complexity, Vocabulary Complexity
General Communication Test Readability, Harmfulness, Sentence Complexity - 5.4k
Vocabulary Complexity, Grammaticality
General Visual Conv Test Fluency, Coherence, Grammaticality, Readability - 1.6k
Text-rich Und Test Fluency, Coherence, Grammaticality, Readability - 1.6k
Robustness-oriented Instructions Test Readability, Bias, Fluency, Coherence - 1.6k
Medical Image Und Test Fluency, Coherence, Readability, Grammaticality - 1.6k
Simple Image Captioning Test Fluency, Clarity, Ethnic Bias, Coherence, - 2.4k
Sentence Complexity, Vocabulary Complexity
Detailed Image Captioning Test Readability, Grammaticality - 1.6k
Sentence Complexity, Vocabulary Complexity
Visual Story Completion Test Readability, Coherence, Grammaticality - 3.8k
Image Repetition, Semantic Consistency
Multimodal Script Generation Test Grammaticality, Fluency - 2.4k
Scene Consistency, Clarity, Image Coherence
FRA-OOD Question Generation Test Clarity, Grammaticality, Fluency, Coherence Engagement, Difficulty 1.1k
Bias, Structure Accuracy, Spelling Accuracy
Title Generation Test Readability, Clarity, Bias, Ethical Bias Suitability, Coverage, Appeal 0.4k
Coherence, Fluency, Grammaticality, Cultural Bias
Keywords Extraction Test Readability, Grammaticality Relevance, Representation 0.5k
Clarity, Fluency, Coherence
Data Analysis Test Coherence, Privacy Violations, Readability, Fluency Precision, Format 1.5k
Clarity, Vocabulary Complexity, Sentence Complexity
Translation Test Structure Accuracy, Spelling Accuracy, Tense Consistency Translation Fidelity, Thoroughness 1.1k
Fluency, Grammaticality, Clarity, Coherence
Academic VQA Test Fluency, Readability, Vocabulary Complexity Expertise, Correctness, Explainability 3.6k
Sentence Complexity, Coherence, Grammaticality
Medical VQA Test Sentence Complexity, Readability, Vocabulary Complexity Terminology, Transparency 3.6k
Fluency, Coherence, Grammaticality, Clarity
Embodied Decision-makin Test Coherence, Clarity, Grammaticality, Gender Bias Interpretability, Completion, Riskiness 4.8k
Harmfulness, Bias, Readability, Regional Bisa, Toxicity
Text-to-Image Generation (GenAI-Bench) Test Harmfulness, Fidelity Count Alignment, Object Relationship 6.7k
Scene Alignment, Negation, Universal
Object Alignment, Spatial Alignment, Alignment
Action Alignment, Color Alignment, Shape Alignment
Texture Alignment, Attribute Alignment, Comparison
Activitynet Continuation Test Readability, Harmfulness, Coherence, Semantic Consistency Reasonableness, Fulfillment 0.5k
Stylistic Consistency, Action Consistency, Fluency, Clarity
Scene Consistency, Image Repetition, Character Consistency
Storytelling Generation Test Readability, Harmfulness, Scene Consistency, Image Coherence Originality, Simplification 0.6k
Stylistic Consistency, Action Consistency, Semantic Consistency
Image Repetition, Character Consistency, Fluency, Coherence
Table 12: Aspects and sample counts in the Training Set, FRA-ID, FRAUAs-OOD, and FRA-OOD. Blue-marked aspects are human-annotated, while the rest are annotated by GPT-4o.
Aspect Abbr Source Target Definition
Bias Bias  [44, 45, 46, 45] T, ITI Bias evaluates the presence and extent of gender, regional, and cultural bias in generated content. It examines whether the content reinforces stereotypes, shows favoritism or discrimination toward specific regions, or reflects racial or ethnic prejudice. The assessment focuses on language, tone, and representation, aiming to identify harmful biases or imbalanced portrayals.
Gender Bias GenB  [45, 47, 48, 48] T, ITI This criterion evaluates whether the generated content reinforces traditional gender stereotypes or presents biased views based on gender. It examines whether certain roles, traits, or behaviors are unfairly associated with specific genders, such as linking nurturing to women or leadership to men.
Cultural Bias CulB  [45] T, ITI This criterion evaluates the model’s understanding of and respect for cultural diversity. It assesses whether the generated text appropriately acknowledges cultural differences and avoids reinforcing harmful generalizations or stereotypes about specific cultures. For instance, models may misinterpret culturally specific idioms or expressions, or fail to recognize regional language variants, which can result in outputs that lack cultural sensitivity or misrepresent certain communities.
Regional Bias RegB  [45, 48] T, ITI This criterion evaluates whether the generated content shows a bias toward or against any particular nationality or region. It ensures that all national or regional identities are treated fairly, without favoritism, discrimination, or negative stereotyping.
Ethnic Bias EthB  [49, 47, 48] T, ITI This criterion assesses whether the generated content reflects racial or ethnic biases, either by perpetuating stereotypes or marginalizing certain groups. It looks for whether certain ethnic or racial groups are portrayed negatively or unfairly advantaged, reinforcing harmful biases.
Toxicity Tox  [50, 46, 51] T, ITI This criterion evaluates whether the model avoids generating harmful, violent, explicit, or inappropriate content. It ensures the system does not produce material that could cause emotional harm, violate legal norms, or infringe on privacy. This includes preventing the generation of discriminatory, NSFW, or harmful language and images.
Harmfulness Har  [52, 46, 51] T, ITI This criterion evaluates whether the generated content promotes, glorifies, or depicts any form of violence, abuse, or harm—be it physical, emotional, or psychological. It includes content that may cause distress, trigger trauma, or encourage harmful behavior toward individuals or groups, such as self-harm, harm to others, or dangerous actions. Images that cause psychological discomfort when seen are harmful.
Privacy Violation Priv  [53, 46, 51] T, ITI This criterion evaluates whether the generated content reveals personal, private, or sensitive information about individuals without consent, violating privacy norms or laws. This includes disclosing private details such as names, locations, health conditions, or other confidential information. The content should respect an individual’s privacy rights, in accordance with legal and ethical standards.
Readability Rea  [54, 14, 8, 55] T, ITI Readability measures how easily a text is understood by its intended audience. It focuses on overall fluency and the clarity of ideas to ensure they are presented simply and align with the readers’ comprehension levels. It evaluates the quality of both inter-sentences and intra-sentences, checking if they are grammatically correct, naturally written, with clear meanings, and good context-relatedness and coherence, allowing readers to absorb information effortlessly.
Coherence Coh  [5, 14, 8] T, ITI Coherence refers to the logical consistency and interconnection of ideas within the text. It ensures that the text follows a clear, organized progression, with each section logically linking to the next. A coherent text allows the reader to easily follow the author’s argument or narrative, with smooth transitions and no sudden shifts in topic or gaps in reasoning.
Clarity Cla  [8, 38] T, ITI Clarity measures how easily the text can be understood. It emphasizes the use of straightforward, unambiguous language, avoiding complex vocabulary and convoluted sentence structures that could confuse the reader.
Text Repetition TexR  [8, 56] T, ITI Text repetition refers to the unnecessary duplication of phrases, sentences, or ideas that do not add new meaning. Repetition disrupts the flow, reduces clarity, and can make the text feel redundant.
Sentence Complexity SenC  [57, 58] T, ITI This criterion evaluates how sentence structures affect readability. Complex or overly intricate sentences may hinder comprehension, while simpler structures enhance clarity.
Table 13: The definition and source of Universal Aspects. (Part 1)
Aspect Abbr Source Target Definition
Vocabulary Complexity VocC  [57, 58] T, ITI This criterion evaluates whether the vocabulary used in the text is well-suited to the target audience’s comprehension level. The language should prioritize clarity by avoiding overly complex or obscure terminology that may hinder understanding, while also ensuring it is not excessively simplistic for the intended readers.
Fluency Flu  [8, 38, 5, 14] T, ITI Fluency measures the smoothness of the language and sentence construction, ensuring the text flows seamlessly and is grammatically correct. The fluent text flows naturally, uses familiar phrasing, and avoids awkward or forced transitions.
Grammaticality Gra  [8, 14] T, ITI Grammaticality measures how well the text adheres to the rules of grammar, including sentence structure, verb tense, subject-verb agreement, and punctuation.
Spelling Accuracy SpeA  [8, 14] T, ITI This criterion assesses the correctness of spelling in the text, ensuring there are no typographical errors or misspelled words.
Tense Consistency TenC  [8, 14] T, ITI This criterion assesses whether the sentence maintains consistent tense usage (past, present, future) throughout. The tense should remain uniform unless a change is contextually required, such as when indicating a change in the time frame.
Structure Accuracy StrA  [8, 14] T, ITI This criterion evaluates the grammatical accuracy of sentences, focusing on the proper alignment of subject-verb-object and the correct linkage of main and subordinate clauses. It ensures that sentences are constructed following standard grammatical rules, which enhances clarity and readability.
Image Quality ImaQ  [59, 9, 10] I, ITI Image quality primarily evaluates the fidelity of the generated image, ensuring it accurately represents the described objects and their characteristics. It also assesses whether the image contains any harmful elements, such as offensive content, misleading representations, or harmful stereotypes.
Fidelity Fid  [9, 10] I, ITI Fidelity evaluates how accurately the generated image represents the shape, characteristics, and behavior of the objects described, ensuring they align with real-world expectations. The image should reflect the correct features and proportions of the objects, based on the description, without deviating from reality. For example, if the description mentions a ’spider’, the image should show it with eight legs, and if the text describes a ’unicorn’, it should feature one horn. Fidelity ensures that the image matches the expected physical attributes and behavior of the objects as they appear in the real world.
Aesthetic Aes  [60, 10] I, ITI Aesthetic measures how pleasing and emotionally striking an image is. It checks if the image combines elements like color and light to create a visually appealing experience that resonates with viewers. This measure helps assess whether the image captures beauty and evokes an emotional response, making it memorable and impactful.
Image Coherence ImaC  [59] ITI Image coherence measures the consistency in style and subject representation across images. This includes textures, color palettes, lighting, rendering styles, and maintaining consistent physical attributes, clothing, and behavioral traits. Image coherence also penalizes image duplication, where the output images are too similar, or within the output images themselves.
Image Repetition ImgR  [59] ITI This criterion penalizes the repetition or excessive similarity between images. It identifies cases where multiple images are too similar, either due to identical content, framing, or visual elements, without adding meaningful variety to the overall set.
Stylistic Consistency StyC  [59] ITI This criterion evaluates the uniformity of the artistic style across different images. It requires that all images share the same visual style, such as all realistic photographs, all cartoon illustrations, or all oil paintings. For example, if one image is a cartoon-style illustration, all related images should follow the same cartoon style, rather than mixing different styles like realism, oil painting, or animation.
Semantic Consistency SemC  [59, 61, 62] ITI This criterion evaluates the consistency of semantic elements in the images across a series. These semantic elements include object attributes, scenes, etc. It ensures that each element within the sequence aligns with the overarching theme or purpose, promoting a clear and unified narrative or instructional flow. For example, a cooking tutorial showing step-by-step cake preparation maintains coherence by following the logical sequence of the recipe, whereas mixing unrelated steps disrupts the flow and confuses the viewer.
Table 14: The definition and source of Universal Aspects. (Part 2)
Aspect Abbr Source Target Definition
Character Consistency ChaC  [59, 61, 62] ITI This criterion ensures that objects or characters maintain the same attributes (e.g., color, shape, size) across different images. It checks that the physical characteristics of a character or object remain constant. This prevents inconsistencies, such as a character suddenly wearing a different outfit or an object changing in size or appearance without explanation.
Action Consistency ActC  [59, 61, 62] ITI Action Consistency in image sequences ensures that actions depicted across multiple images follow a logical progression and match in activity and intent. This helps maintain a smooth narrative flow, such as a person going from jogging to running, rather than abruptly switching to an unrelated action like swimming.
Scene Consistency SceC  [61, 62] ITI This criterion evaluates how effectively a sequence of images maintains the consistency of the scene elements, such as location and background. For instance, in a photo story about a day at the farmer’sbenchmarket, each image should consistently depict thebenchmarket environment from dawn to dusk. If a photo suddenly shifts to a beach setting, it disrupts the scene consistency, confusing the viewer and breaking the narrative flow.
Text-Image Relationship T-IR  [59] ITI Text-Image Relationship refers to how text and image work together to enhance each other. This includes how the text describes what’s shown in the image—objects, their properties, and spatial relationships—as well as providing extra context or background information to add meaning beyond mere description. Text doesn’t always need to explain the image directly; it can also convey emotions or offer complementary details to deepen understanding. Effective text-image alignment ensures the text not only describes the image but also enriches its meaning.
Alignment Ali  [11, 16, 15, 9] ITI Alignment between text and image is evaluated based on how well the image matches the details in the text. This includes the correct objects, their attributes (e.g., color, size), spatial relationships (e.g., left, right), and actions. The alignment should consider multiple dimensions. For example, if the text describes “two boys wearing white clothes, one on the left and one on the right, playing badminton,” the image should show two boys in white clothes, placed correctly in the scene, with a badminton game happening between them.
Object Relationship ObjR  [25, 16] ITI Object Relationship evaluates how accurately the spatial relationships and actions between objects described in the text are represented in the image. It focuses on whether the interactions, positions, and movements of objects in the image match those described in the text. For example, in the description ’a person standing next to a table and reaching for a book’, the image should show the person next to the table with their hand extended toward the book. As long as the spatial relationships and actions are correctly represented, the image will be considered accurate.
Spatial Alignment SpaA  [63, 15, 25, 16] ITI This criterion measures how accurately the spatial relationships between objects in the generated output match the descriptions in the text. It evaluates whether the relative positions, orientations (up, down, left, right), and arrangements of objects align with the textual description. For example, if the text mentions a chair to the left of a table, the image should show the chair on the left side of the table, and if a person is sitting on a chair in front of a table, this relationship should also be reflected accurately in the image.
Action Alignment ActA  [25, 16] ITI Action Alignment evaluates how accurately the actions or interactions described in the text are represented in the image. It specifically measures dynamic actions like ’running’, ’talking’, or ’holding’, and does not consider static descriptions or spatial relationships (e.g., left, right, or object arrangements). For example, if the text says ’a dog running after a ball’, the image should show the dog running after the ball.
Count Alignment CouA  [25, 16, 15] ITI Count Alignment evaluates how accurately the number of objects described in the text matches both the number and type of objects presented in the corresponding image. It specifically measures whether the quantity of objects mentioned in the text (e.g., two dogs, three cars, or five apples) is correctly reflected in the image, and ensures that the objects shown are of the correct type. For example, if the text mentions ’two dogs’, the image should contain two dogs, not two cats. This evaluation focuses solely on the count and type of objects, without considering their attributes or positions.
Object Alignment ObjA  [25, 15] ITI Object Alignment evaluates how accurately the objects described in a text are represented in an image. The objects mentioned in the text should be accurately presented in the image, with no omissions. For example, in the description ’ A man sitting by himself at the dinner table’, the image should feature both a man and a dinner table to achieve a high score. Missing or extra objects in the image would lower the score.
Table 15: The definition and source of Universal Aspects. (Part 3)
Aspect Abbr Source Target Definition
Scene Alignment ScnA  [25] ITI Scene Alignment assesses the extent to which the setting or background described in a text is accurately represented in an image. It focuses on elements such as location, weather conditions, time of day, and general atmosphere rather than specific objects. For example, if a text describes ’a sunny environment with historical architecture around a town square’, the image should depict this specific setting, showing historical architecture situated within a town square context, not isolated in an open field or desert.
Attribute Alignment AttA  [25, 16, 15] ITI Attribute Alignment assesses how well the attributes (such as color, texture, size, and shape) of objects described in the text are reflected in the image. It ensures that key features like color and size match what’s described. For instance, if the text says “a green bicycle,” the image should show a bicycle that is green. It also checks if details like texture and shape are accurate, like a “smooth, red ball” being depicted as both smooth and red. Furthermore, the relative size and proportions of objects should be consistent with the description. Overall, the image should visually match the text’s description of each object’s attributes.
Color Alignment ColA  [25, 16, 15] ITI This criterion assesses whether the colors described in the text match the corresponding objects in the image based solely on the text description, or whether the colors of the objects conform to realistic colors. If the text specifies the color of an object, the evaluation checks whether the image reflects that color. For example, if the text says ’a blue car’, the image should show a car that matches the described color, not one of a different color.
Shape Alignment ShaA  [25, 16, 15] ITI This criterion evaluates how accurately the shapes described in the text correspond to the shapes of objects in the image, based solely on the text’s shape descriptors, or whether the shape of the objects conforms to realistic shape. The text may include common shape terms such as square, round, rectangular, oval, triangular, cubic, cylindrical, spherical, and other basic or specialized shapes like pyramidal, conical, pentagonal, teardrop, crescent, and diamond.
Texture Alignment TexA  [25, 16, 15] ITI This criterion evaluates how accurately the textures described in the text are reflected in the image, based solely on the text description, or whether the textures of the objects conform to realistic textures. It includes surface qualities such as smoothness, roughness, or granularity, as well as material types like wood, plastic, and rubber. Additionally, it may encompass tactile sensations or patterns, such as silky, polished, or other material-based descriptions. The accuracy of the texture description is assessed by comparing the textual description to the actual appearance and surface properties of the objects in the image.
Advanced Alignment AdvA  [25] ITI Advanced Alignment is a comprehensive evaluation aspect that assesses the ability to accurately represent complex textual descriptions in an image. It encapsulates Comparison, Negation, and Universality aspects. This means it checks whether the image correctly depicts comparisons between entities based on attributes such as number or volume; represents absence or contradiction of elements indicated by words like ’no’ or ’not’; and illustrates scenarios where every member of a group shares a specific attribute or relation.
Comparison Cmpa  [25] ITI Comparison evaluates the ability to accurately depict comparisons between entities as described in a text. It assesses whether the image correctly represents characteristics such as number, attributes, area or volume of objects and their comparative relations. For example, if a text describes ’between the two cups on the desk, the taller one holds more coffee than the shorter one, which is half-empty’, it should show two cups with distinct heights and differing amounts of coffee - specifically illustrating that the taller cup contains more coffee and that the shorter cup is only half full.
Negation Nega  [25] ITI Negation evaluates the ability to accurately depict the absence or contradiction of elements as described in a text. It assesses whether the image correctly represents scenarios where entities are not present, as indicated by words such as ’no’, ’not’, or ’without’. For example, if a text describes ’a bookshelf with no books, only picture frames’, it should show a bookshelf devoid of books but filled with picture frames instead.
Table 16: The definition and source of Universal Aspects. (Part 4)
Aspect Abbr Source Target Definition
Universal Univ  [25] ITI Universal evaluates the ability to accurately depict scenarios where every member of a group shares a specific attribute or is involved in a common relation, as described in a text. It assesses whether the image correctly represents situations indicated by words such as ’every’, ’all’, ’each’, or ’both’. For example, if a text describes ’a bustling kitchen where every chef is preparing a dish’, it should show all chefs within the kitchen actively engaged in cooking.
Complement Comp  [64] ITI This criterion assesses how text and image work together to enhance or clarify the overall meaning. It evaluates how they complement each other, whether through redundancy, emotional amplification, added context, or interpretation, creating a richer understanding than either could provide alone.
Complementary Overlap ComO  [64] ITI This criterion captures cases where the image and text redundantly represent the same information, and the image does not contribute anything beyond what is already stated in the text. The text alone is sufficient to convey the core message. For example, a caption saying “Tacos are the best” can be paired with a photo of tacos. Both the image and text communicate the same idea with no additional contextual or semantic contribution from the image.
Emotional Complement EmoC  [64] ITI This criterion evaluates how the image, particularly through elements like emojis or memes, enhances or clarifies the emotional tone of the text. The image should convey emotions, moods, or subtle feelings that the text alone might not fully express. This could involve using visual elements such as facial expressions, symbols, or humorous images that directly amplify, complement, or interpret the emotional content of the text, helping to create a more vivid or relatable emotional context.
Comment Complement ComC  [64] ITI This criterion assesses how the text offers commentary or evaluation that adds meaning to the image. The text may provide an interpretation, critique, or personal perspective on the image, offering a deeper understanding or a subjective view of what the image represents. The text can either affirm or challenge the implied meaning of the image, creating a more nuanced interaction between the two modalities.
Contextual Complement ConC  [64] ITI This criterion evaluates how the text and the image provide complementary background information or contextual details that enhance the understanding of each other. The text should add context, explanations, or relevant background information that helps interpret the image more fully, or conversely, the image should provide visual context that clarifies or enriches the meaning of the text.
Table 17: The definition and source of Universal Aspects. (Part 5)
Aspect Abbr Source Sub-Task Definition
Coverage Cov  [38] Summarization The summary should cover all essential points from the original text, ensuring no significant information is omitted. In this case, completeness is more important than brevity.
Length Constraint LenC  [38] Summarization This measures the model’s ability to respect any given length restrictions for the response (e.g., sentence count, paragraph count, or specific character limits), ensuring the model does not generate an overly short or excessively long response. If the instruction does not mention a length requirement, all responses should receive a score of -1, without considering the influence of other criteria.
Layout Lay  [38] Summarization The written summary is encouraged to follow a clear and well-organized layout. You may use headings, bullet points, lists, tables, or other formatting devices to improve readability. This criterion is not applicable if the user does not explicitly require a specific layout.
Conciseness Con  [38] Summarization This criterion evaluates whether the summary is brief and to the point, effectively capturing the essential information of the original content while avoiding unnecessary details or excessive elaboration. It focuses on the ability to convey key elements clearly and concisely, maintaining the core meaning of the text without including irrelevant or redundant content.
Table 18: The definition and source of Task-specific Aspects. (Part 1)
Aspect Abbr Source Sub-Task Definition
Accuracy Acc  [38] Rewriting Accuracy in Rewriting refers to how correctly the changes follow the given instructions. Any modifications are made according to the task requirements.
Pointing Out PoiO  [38] Rewriting This criterion evaluates the clarity and transparency of corrections made to the text. Instead of simply returning a corrected version, the correction process must be clearly outlined. This includes pointing out which parts of the text were changed, explaining how the changes were made, and providing a reason for each correction. This approach helps the user understand not only the mistake but also the logic behind the correction, facilitating better learning and comprehension.
Difference Dif  [38] Rewriting This criterion evaluates whether the written text shows a clear difference from the original in terms of its appearance or structure. The text should not simply be a direct copy of the original; instead, it should reflect meaningful changes, such as rephrasing or reformatting. The goal is to ensure that the new text is distinct and provides added value, rather than just repeating the same content in the same form.
Creativity Cre  [38] Creative Writing Creativity refers to the originality and imagination in the writing. It evaluates whether the content presents fresh ideas, unique expressions, and an inventive approach. Creativity involves not only the ideas but also how the language, structure, and emotions are used. It looks for originality that breaks from the norm, offering something new and engaging for the reader.
Engagingness Eng  [38] Creative Writing Engagingness refers to the extent to which a text, whether in creative writing, captures and maintains the audience’s interest. This criterion evaluates the ability of the content to stimulate a compelling exchange of ideas, emotions, or information, encouraging active participation and sustained engagement. It measures how well the narrative or dialogue keeps the audience engaged and thought-provoking throughout.
Theme The  [38] Creative Writing This criterion evaluates whether the response maintains a clear and consistent central theme or subject matter throughout the creative work, such as a song or piece of writing. For example, if the song centers around a romantic breakup, every verse and chorus should consistently reflect and support this theme, creating a cohesive narrative. The goal is to ensure that all elements of the work work together to reinforce the main idea, helping the audience to easily follow and connect with the message being conveyed.
Relevance Rel  [38] Functional Writing Relevance measures how closely the content of a written piece aligns with the specific objectives and purpose of the task. It ensures that the text stays on topic, directly addresses the intended subject, and provides information that is appropriate and meaningful for the given context.
Contextual Information ConI  [38] Functional Writing This criterion assesses how effectively extra context or elaboration is provided when necessary, enhancing the user’s understanding without overwhelming them. It ensures that the information is valuable and directly aligned with the user’s inquiry, offering clarity and depth where needed
Example Quality ExaQ  [38] Functional Writing This criterion assesses how effectively specific instances or cases are provided to illustrate the concept being discussed. Examples help make abstract ideas more tangible and understandable by showing how they manifest in real-world or hypothetical scenarios.
Accuracy Acc  [38] General Communication Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
Information Richness InfoR  [38] General Communication This criterion assesses whether the response provides sufficient information to address the query, including necessary context and explanations. It ensures that the answer is detailed and clear, offering enough depth to fully address the user’s needs without omitting key information.
Understandability Und  [38] General Communication This criterion evaluates understandability, which refers to how well a conversational system interprets user inputs and generates contextually relevant responses. It focuses on the model’s ability to grasp the meaning of user queries and the flow of the conversation, ensuring responses are appropriate. For example, if a user asks, ’What’s the weather like today?’, the model should respond with a relevant, clear answer about the current weather, instead of a generic or unrelated reply. A model with high understandability maintains seamless, clear communication.
Table 19: The definition and source of Task-specific Aspects. (Part 2)
Aspect Abbr Source Sub-Task Definition
Suitability Suit  [38] Title Generation This criterion evaluates how closely the generated title aligns with the content of the text or work. It’s about ensuring that the title directly reflects the main theme or subject matter, providing a meaningful and appropriate summary for the given context. For instance, if the text is about the effects of climate change on wildlife, a relevant title might be Wildlife in Distress: Unraveling the Impact of Climate Change.
Coverage Cov  [38] Title Generation This criterion evaluates how much of the essential information from the source content is captured in the generated title. It’s about assessing how well the title encapsulates the key points or main ideas of the source content. For instance, if the text discusses the history, types, and health benefits of meditation, a title that covers these points might be Meditation: Tracing Its Roots and Understanding Its Health Wonders.
Appeal Appe  [38] Title Generation This criterion evaluates the attractiveness of the generated title. It assesses whether the title has the power to catch the reader’s attention, spark interest, and entice the reader to delve into the content. For example, a title like ’Unlocking Creativity: Harnessing the Power of an Unconventional Mind’ might appeal to readers interested in personal development and creativity.
Translation Fidelity TrnF  [38] Translation This criterion evaluates the precision of the translation in replicating the source text’s content in the target language. It involves correct vocabulary choices, proper grammar, and syntax. It ensures that the meaning, tone, and nuances of the original text are preserved without adding, omitting, or distorting information.
Thoroughness Thor  [38] Translation This criterion assesses whether the translation covers all parts of the source text without leaving out any information, details, or nuances. It ensures that the translation fully conveys the original text’s content, including all its ideas, arguments, and points, in the target language.
Precision Prec  [38] Data Analysis This criterion evaluates the precision and correctness of the data analysis results. It assesses the degree to which the results correspond to the true or actual values, ensuring that the output is free from errors or distortions.
Format Form  [38] Data Analysis This criterion evaluates the clarity and user-friendliness of the data analysis output generated by the model. It assesses whether the output is well-structured, follows standard data presentation conventions, and is easy to interpret and understand. The output should present the analysis results in a clear and concise manner, using appropriate tables, charts, or other visual aids as necessary. The use of clear headings, labels, and explanations should also be considered.
Engagingness Eng  [38] Question Generation This criterion evaluates the ability of the generated questions to captivate and hold the attention of the audience. Engaging questions spark curiosity, stimulate thought, and encourage active participation. For instance, a question like ’How does climate change impact our daily lives?’ might engage readers because it connects a global issue to personal experiences.
Difficulty Diff  [38] Question Generation This criterion assesses whether the complexity level of the generated questions matches the knowledge and comprehension level of the intended audience. For example, for a high school audience, a question like ’Can you explain Einstein’s theory of relativity?’ might be too difficult, while ’What is the capital of France?’ might be too easy.
Relevance Rel  [38] Keywords Extraction In the context of keyword extraction, this criterion evaluates how closely the extracted keywords align with the main themes or subjects of the text. Relevant keywords should accurately reflect the core ideas, topics, or concepts presented in the text, providing a concise and meaningful representation of the content.
Representation Repr  [38] Keywords Extraction In the context of keyword extraction, this criterion assesses the extent to which the extracted keywords represent all the significant themes, topics, or concepts in the text. A high coverage means that the keywords collectively provide a comprehensive summary of the text, leaving no important aspect unrepresented.
Helpfulness Hel  [65, 66] General Visual Conversation Consider whether the generated text provides valuable insights, additional context, or relevant information that contributes positively to the user’s comprehension of the image. Assess whether the model accurately follows any specific instructions or guidelines provided in the prompt. Evaluate the overall contribution of the response to the user experience.
Instruction Following InsF  [18, 66] General Visual Conversation Instruction Following is defined as the ability to accurately interpret and execute the task as outlined in the given instruction, focusing solely on providing the correct answer. It measures whether the response directly addresses the task’s core requirements, without considering additional background information, explanations, tone, or length. Even if a single word answers the question correctly, the response should receive a high score on this criterion.
Table 20: The definition and source of Task-specific Aspects. (Part 3)
Aspect Abbr Source Sub-Task Definition
Accuracy Acc  [65, 67] Text Rich Understanding Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
Instruction Following InsF  [65, 18] Text Rich Understanding Instruction Following is defined as the ability to accurately interpret and execute the task as outlined in the given instruction, focusing solely on providing the correct answer. It measures whether the response directly addresses the task’s core requirements, without considering additional background information, explanations, tone, or length. Even if a single word answers the question correctly, the response should receive a high score on this criterion.
Instruction Inconsistency InsI  [65, 19] Robustness-oriented Instructions Instruction inconsistency refers to the model’s outputs that deviate from a user’s directive. While some deviations might serve safety guidelines, the inconsistencies here signify unintentional misalignment with non-malicious user instructions. For example, the user’s actual intention is translation, However, the LLM erroneously deviated from the user’s instruction and performed a question-answering task instead.
Context Inconsistency ConI  [65, 19] Robustness-oriented Instructions Context inconsistency occurs when the model’s output contradicts the contextual information provided by the user, which may include text or images. This happens when the model misinterprets the context or incorrectly identifies elements, such as attributes of objects in an image or non-existent details. For example, if the user states that the source of the Nile is in the Great Lakes region of central Africa, but the model incorrectly claims it is in the Americas. Similarly, if an image shows a girl in red clothes, but the model mistakenly identifies her as a boy in blue clothes.
Accuracy Acc  [68, 65] Medical Image Understanding Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
Instruction Following InsF  [18, 69] Chart Reasoning Instruction Following is defined as the ability to accurately interpret and execute the task as outlined in the given instruction, focusing solely on providing the correct answer. It measures whether the response directly addresses the task’s core requirements, without considering additional background information, explanations, tone, or length. Even if a single word answers the question correctly, the response should receive a high score on this criterion.
Accuracy Acc  [65, 69] Chart Reasoning Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
Instruction Following InsF  [18, 70] Graph Reasoning Instruction Following is defined as the ability to accurately interpret and execute the task as outlined in the given instruction, focusing solely on providing the correct answer. It measures whether the response directly addresses the task’s core requirements, without considering additional background information, explanations, tone, or length. Even if a single word answers the question correctly, the response should receive a high score on this criterion.
Accuracy Acc  [65, 70] Graph Reasoning Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
Instruction Following InsF  [18, 71] Text Reasoning Instruction Following is defined as the ability to accurately interpret and execute the task as outlined in the given instruction, focusing solely on providing the correct answer. It measures whether the response directly addresses the task’s core requirements, without considering additional background information, explanations, tone, or length. Even if a single word answers the question correctly, the response should receive a high score on this criterion.
Accuracy Acc  [65, 71] Text Reasoning Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
Table 21: The definition and source of Task-specific Aspects. (Part 4)
Aspect Abbr Source Sub-Task Definition
Conciseness Con  [72, 73] Simple Image Captioning This criterion examines whether the caption remains concise and focused, conveying the necessary information efficiently without including redundant or overly detailed content. It emphasizes clarity and precision in communicating key points while minimizing distractions from irrelevant elements.
Object Alignment ObjA  [72, 73] Simple Image Captioning Object Alignment evaluates the precision with which images match the objects described in the corresponding text. This criterion ensures that the objects mentioned in the text are accurately represented in the image, focusing solely on whether the object itself is present. It does not consider additional criteria such as attribute alignment.
Text-Image Relationship T-IR  [72, 73] Simple Image Captioning Text-Image Relationship refers to how text and image work together to enhance each other. This includes how the text describes what’s shown in the image—objects, their properties, and spatial relationships—as well as providing extra context or background information to add meaning beyond mere description. Text doesn’t always need to explain the image directly; it can also convey emotions or offer complementary details to deepen understanding. Effective text-image alignment ensures the text not only describes the image but also enriches its meaning.
Alignment Ali  [72, 73] Simple Image Captioning Alignment between text and images ensures the image matches the description, including the attributes and relationships of objects. For example, if the text is ’a cat dressed as Napoleon Bonaparte,’ the image must show a cat. If the text is ’a little girl sitting in front of a sewing machine,’ the image should show only one girl in that setting, not multiple girls. Misalignment happens when the subject is missing or the image doesn’t match the description.
Completeness Com  [73, 74] Detailed Image Captioning Completeness evaluates whether the caption covers all significant elements in the image, including key objects, actions, and possible contextual information (e.g., background, setting, or scene details). A complete caption should mention all essential features of the image without omitting critical information.
Attribute Alignment AttA  [73, 74] Detailed Image Captioning Attribute Alignment assesses how well the attributes (such as color, texture, size, and shape) of objects described in the text are reflected in the image. It ensures that key features like color and size match what’s described. For instance, if the text says “a green bicycle,” the image should show a bicycle that is green. It also checks if details like texture and shape are accurate, like a “smooth, red ball” being depicted as both smooth and red. Furthermore, the relative size and proportions of objects should be consistent with the description. Overall, the image should visually match the text’s description of each object’s attributes.
Spatial Alignment SpaA  [73, 74] Detailed Image Captioning In the image captioning generation task, this criterion measures how accurately the spatial relationships between objects in the text align with their representation in the image. It evaluates whether the relative positions and arrangements of objects described in the text are correctly reflected in the image, focusing solely on spatial relationships. It does not consider other criteria, such as object attributes or additional context.
Count Alignment CouA  [73, 74] Detailed Image Captioning Count Alignment evaluates how accurately the number of objects described in the text matches both the number and type of objects presented in the corresponding image. It specifically measures whether the quantity of objects mentioned in the text (e.g., two dogs, three cars, or five apples) is correctly reflected in the image, and ensures that the objects shown are of the correct type. For example, if the text mentions ’two dogs’, the image should contain two dogs, not two cats. This evaluation focuses solely on the count and type of objects, without considering their attributes or positions.
Action Alignment ActA  [73, 74] Detailed Image Captioning In the image captioning generation task, Action Alignment evaluates how accurately the actions or interactions depicted in the image are described in the text. It specifically measures actions like ’running’, ’talking’, or ’holding’, and does not take into account static descriptions or spatial relationships (e.g., left, right, or object arrangements). For example, if the image shows a dog running after a ball, the text should include the word ’running’.
Object Alignment ObjA  [73, 74] Detailed Image Captioning Object Alignment evaluates the precision with which images match the objects described in the corresponding text. This criterion ensures that the objects mentioned in the text are accurately represented in the image, focusing solely on whether the object itself is present. It does not consider additional criteria such as attribute alignment.
Table 22: The definition and source of Task-specific Aspects. (Part 5)
Aspect Abbr Source Sub-Task Definition
Color Alignment ColA  [73, 74] Detailed Image Captioning In the image captioning generation task, this metric assesses whether the colors described in the text match the corresponding objects in the image. If the image presents an object with a specific color, the text should describe it as accurately as possible. For example, if the image shows a blue car and a person wearing a yellow shirt, the text should mention ’blue’ and ’yellow’. The more accurately the colors are described, the higher the score.
Texture Alignment TexA  [73, 74] Detailed Image Captioning In the image captioning generation task, this metric checks how accurately the textures presented in the image are described in the text. If the objects in the image have specific texture qualities, such as surface qualities (e.g., smooth, rough), material types (e.g., wooden, metallic), or tactile sensations or patterns (e.g., silky, polished), the more accurately these textures are expressed in the text, the higher the score.
Shape Alignment ShaA  [73, 74] Detailed Image Captioning In the image captioning generation task, this metric assesses how accurately the shapes described in the text match the shapes of objects in the image. If the image contains objects with shape descriptors such as geometric terms (e.g., cubic, circular, triangular) or more general attributes like symmetry, curvature, complexity, and flatness, the text should include corresponding descriptions. The more accurately the shapes are described, the higher the score.
Expertise Expe  [75] Academic QA Expertise refers to the model’s ability to accurately use domain-specific terminology when answering academic questions. This includes correctly applying terms relevant to the subject, such as those in science, history, or literature. For example, when asked about a historical event, the model should use terms like ’primary sources’ or ’historical context’ instead of vague words like ’thin’ or ’event’. Proper use of academic terms shows the model’s understanding of the subject.
Correctness Corr  [75] Academic VQA The conciseness metric assesses whether the model provides clear, direct, and succinct answers to academic questions, avoiding unnecessary details. The accuracy metric evaluates if the responses correctly address the query based on the provided academic context, ensuring precise and relevant information is delivered.
Explainability Expl  [75] Academic VQA Explainability is a criterion used to assess how clearly a model explains its reasoning when answering an academic question. It focuses on whether the model provides clear, logical steps and justifications for its answer, making it easy for users to understand how the answer was reached.
Interpretability Intp  [76] Embodied Decision-making Interpretability in the context of embodied decision-making refers to how well a model can articulate its decisions and actions. It’s about transparency, where the model should be able to provide clear explanations for why it chose a particular course of action over others. For example, if an AI playing chess makes a move, interpretability would involve the AI being able to explain why it made that specific move, considering factors like board position, potential threats, and future opportunities.
Completion Comp  [76] Embodied Decision-making Task completion is essentially about goal achievement. In embodied decision-making tasks, this means whether or not the model has successfully navigated its environment to achieve its intended objective. This could range from simple objectives, such as reaching a certain location within an environment (like an AI navigating through a maze), to more complex ones, like completing multiple interrelated tasks in sequence (like cooking, where you need to chop vegetables before frying them).
Riskiness Risk  [76] Embodied Decision-making "Riskiness measures how well the model avoids harmful situations or risky behaviors during task execution. This is particularly important in real-world applications where mistakes can have serious consequences. For instance, if we consider autonomous driving again, safety here doesn’t just mean avoiding collisions; it also involves obeying traffic rules, not taking unnecessary risks while overtaking other vehicles, and maintaining safe distances from pedestrians and other road users.
Terminology Term  [77] Medical VQA Terminology evaluates the model’s ability to correctly use and apply medical terminology when responding to queries related to medical image understanding. This includes the accurate usage of terms specific to anatomy, pathology, imaging techniques, or diagnostic processes. For example, if a user asks about a CT scan showing a lung lesion, the model should mention terms like ’pulmonary nodule,’ ’radiographic features,’ or ’lesion characterization’ rather than generic terms like ’spot’ or ’image area.’ Proper use of medical terms indicates the model’s depth of understanding in the field.
Table 23: The definition and source of Task-specific Aspects. (Part 6)
Aspect Abbr Source Sub-Task Definition
Transparency Tran  [77] Medical VQA Transparency is a criterion used to assess how clearly and transparently a model explains its reasoning behind the answer to an academic question. It focuses on the model’s ability to provide understandable, logical, and well-supported justifications for its responses. This includes breaking down complex ideas, referencing relevant information, and demonstrating a step-by-step thought process. A high level of explainability ensures that users not only receive a correct answer but also understand how the answer was derived, helping to build trust in the model’s reasoning and enhancing learning outcomes.
Alignment Ali  [11, 16, 15, 9] Text-to-Image Generation Alignment between text and image is evaluated based on how well the image matches the details in the text. This includes the correct objects, their attributes (e.g., color, size), spatial relationships (e.g., left, right), and actions. The alignment should consider multiple dimensions. For example, if the text describes “two boys wearing white clothes, one on the left and one on the right, playing badminton,” the image should show two boys in white clothes, placed correctly in the scene, with a badminton game happening between them.
Object Relationship ObjR  [25, 16] Text-to-Image Generation Object Relationship evaluates how accurately the spatial relationships and actions between objects described in the text are represented in the image. It focuses on whether the interactions, positions, and movements of objects in the image match those described in the text. For example, in the description ’a person standing next to a table and reaching for a book’, the image should show the person next to the table with their hand extended toward the book. As long as the spatial relationships and actions are correctly represented, the image will be considered accurate.
Spatial Alignment SpaA  [15, 25, 16, 63] Text-to-Image Generation This criterion measures how accurately the spatial relationships between objects in the generated output match the descriptions in the text. It evaluates whether the relative positions, orientations (up, down, left, right), and arrangements of objects align with the textual description. For example, if the text mentions a chair to the left of a table, the image should show the chair on the left side of the table, and if a person is sitting on a chair in front of a table, this relationship should also be reflected accurately in the image.
Action Alignment ActA  [25, 16] Text-to-Image Generation Action Alignment evaluates how accurately the actions or interactions described in the text are represented in the image. It specifically measures dynamic actions like ’running’, ’talking’, or ’holding’, and does not consider static descriptions or spatial relationships (e.g., left, right, or object arrangements). For example, if the text says ’a dog running after a ball’, the image should show the dog running after the ball.
Count Alignment CouA  [25, 16, 15] Text-to-Image Generation Count Alignment evaluates how accurately the number of objects described in the text matches both the number and type of objects presented in the corresponding image. It specifically measures whether the quantity of objects mentioned in the text (e.g., two dogs, three cars, or five apples) is correctly reflected in the image, and ensures that the objects shown are of the correct type. For example, if the text mentions ’two dogs’, the image should contain two dogs, not two cats. This evaluation focuses solely on the count and type of objects, without considering their attributes or positions.
Object Alignment ObjA  [25, 15] Text-to-Image Generation Object Alignment evaluates how accurately the objects described in a text are represented in an image. The objects mentioned in the text should be accurately presented in the image, with no omissions. For example, in the description ’ A man sitting by himself at the dinner table’, the image should feature both a man and a dinner table to achieve a high score. Missing or extra objects in the image would lower the score.
Attribute Alignment AttA  [25, 16, 15] Text-to-Image Generation Attribute Alignment assesses how well the attributes (such as color, texture, size, and shape) of objects described in the text are reflected in the image. It ensures that key features like color and size match what’s described. For instance, if the text says “a green bicycle,” the image should show a bicycle that is green. It also checks if details like texture and shape are accurate, like a “smooth, red ball” being depicted as both smooth and red. Furthermore, the relative size and proportions of objects should be consistent with the description. Overall, the image should visually match the text’s description of each object’s attributes.
Color Alignment ColA  [25, 16, 15] Text-to-Image Generation This criterion assesses whether the colors described in the text match the corresponding objects in the image based solely on the text description, or whether the colors of the objects conform to realistic colors. If the text specifies the color of an object, the evaluation checks whether the image reflects that color. For example, if the text says ’a blue car’, the image should show a car that matches the described color, not one of a different color.
Table 24: The definition and source of Task-specific Aspects. (Part 7)
Aspect Abbr Source Sub-Task Definition
Shape Alignment ShaA  [25, 16, 15] Text-to-Image Generation This criterion evaluates how accurately the shapes described in the text correspond to the shapes of objects in the image, based solely on the text’s shape descriptors, or whether the shape of the objects conforms to realistic shape. The text may include common shape terms such as square, round, rectangular, oval, triangular, cubic, cylindrical, spherical, and other basic or specialized shapes like pyramidal, conical, pentagonal, teardrop, crescent, and diamond.
Texture Alignment TexA  [25, 16, 15] Text-to-Image Generation This criterion evaluates how accurately the textures described in the text are reflected in the image, based solely on the text description, or whether the textures of the objects conform to realistic textures. It includes surface qualities such as smoothness, roughness, or granularity, as well as material types like wood, plastic, and rubber. Additionally, it may encompass tactile sensations or patterns, such as silky, polished, or other material-based descriptions. The accuracy of the texture description is assessed by comparing the textual description to the actual appearance and surface properties of the objects in the image.
Advanced Alignment AdaA  [25] Text-to-Image Generation Advanced Alignment is a comprehensive evaluation aspect that assesses the ability to accurately represent complex textual descriptions in an image. It encapsulates Comparison, Negation, and Universality aspects. This means it checks whether the image correctly depicts comparisons between entities based on attributes such as number or volume; represents absence or contradiction of elements indicated by words like ’no’ or ’not’; and illustrates scenarios where every member of a group shares a specific attribute or relation.
Comparison Cmpa  [25] Text-to-Image Generation Comparison evaluates the ability to accurately depict comparisons between entities as described in a text. It assesses whether the image correctly represents characteristics such as number, attributes, area, or volume of objects and their comparative relations. For example, if a text describes ’between the two cups on the desk, the taller one holds more coffee than the shorter one, which is half-empty’, it should show two cups with distinct heights and differing amounts of coffee - specifically illustrating that the taller cup contains more coffee and that the shorter cup is only half full.
Negation Nega  [25] Text-to-Image Generation Negation evaluates the ability to accurately depict the absence or contradiction of elements as described in a text. It assesses whether the image correctly represents scenarios where entities are not present, as indicated by words such as ’no’, ’not’, or ’without’. For example, if a text describes ’a bookshelf with no books, only picture frames’, it should show a bookshelf devoid of books but filled with picture frames instead.
Universal Univ  [25] Text-to-Image Generation Universal evaluates the ability to accurately depict scenarios where every member of a group shares a specific attribute or is involved in a common relation, as described in a text. It assesses whether the image correctly represents situations indicated by words such as ’every’, ’all’, ’each’, or ’both’. For example, if a text describes ’a bustling kitchen where every chef is preparing a dish’, it should show all chefs within the kitchen actively engaged in cooking.
Accuracy Acc  [78] Image Editing Accuracy measures how well the edited image aligns with the given instructions in the text. The model should accurately implement the specified changes, including adjusting colors, shapes, positions, and adding or removing elements as instructed. This includes adhering to any quantity or specific attribute requirements described in the text. For example, if the text requests ’add one small white dog’, the image should show exactly one small white dog. If the model generates two white dogs or a black dog instead, the accuracy evaluation should be poor.
Creativity Cre  [79, 59] Visual Story Completion Creativity in visual story completion refers to the originality and imagination in developing the narrative. It evaluates how fresh ideas, unique expressions, and an inventive approach are used to expand upon the visual elements. Creativity looks for originality in the interpretation of visuals, offering a new and engaging continuation of the story that captivates the reader.
Engagingness Eng  [79, 59] Visual Story Completion Engagingness refers to the degree to which a story fosters continued interaction, maintains or elevates interest, and stimulates a compelling exchange of ideas, emotions, or information. This criterion assesses the ability of the content to engage the audience, encouraging active participation and sustained engagement by making the narrative or dialogue both interesting and thought-provoking.
Table 25: The definition and source of Task-specific Aspects. (Part 8)
Aspect Abbr Source Sub-Task Definition
Completeness Com  [79, 59] Visual Story Completion Completeness measures whether the model-generated visual story provides a complete narrative, covering all necessary elements (e.g., characters, settings, actions, and events) from the beginning to the end of the story. For example, if the input is a sequence of images showing a person preparing for a hike, the generated story should include all key steps such as packing a backpack, driving to the trail, starting the hike, and reaching the destination, ensuring a coherent and comprehensive narrative.
Completeness Com  [59, 80] Multimodal Script Generation Completeness measures whether the model-generated steps provide a complete guide to finishing the task, covering all necessary actions needed from start to finish.
Feasibility Fea  [59, 80] Multimodal Script Generation This criterion evaluates the feasibility of content output by a model for a specific task by assessing whether each step is factually correct, logically sequenced, and detailed enough for successful execution. It ensures that the steps are not only accurate but also practical, providing clear and actionable instructions that lead to the desired outcome without leading to confusion or errors.
Helpfulness Hel  [59, 80] Multimodal Script Generation This criterion evaluates whether the model’s response provides clear and practical steps to help the user complete the task, such as planting a strawberry. It checks if the response includes all essential details, like preparing the soil, planting the seeds, and caring for the plant, along with useful tips or warnings.
Safety Saf  [59, 80] Multimodal Script Generation This criterion checks if the model’s responses are safe and avoid harmful content. It ensures the model doesn’t suggest dangerous, unethical, or illegal actions, and follows safety and ethical guidelines. The model should also provide safety warnings or advice when needed, prioritizing user well-being and avoiding any risks.
Reasonableness Reas  [59] Storytelling Generation This criterion evaluates the logical coherence and believability of the generated events within the context of the narrative. It assesses whether the events follow a logical progression, align with the established facts and rules of the narrative world, and are consistent with the characters’ motivations and behaviors. The events should not introduce any inconsistencies or contradictions that could disrupt the audience’s suspension of disbelief. Furthermore, the events should be plausible and realistic, avoiding any absurd or improbable scenarios unless they are justified by the narrative context.
Fulfillment Full  [59] Storytelling Generation This criterion evaluates whether the generated events are comprehensive and well-rounded within the narrative. It assesses whether each event has a clear beginning, middle, and end, providing a full and satisfying narrative arc. The events should not feel fragmented or incomplete, but should provide all the necessary details and developments to understand and appreciate their significance in the narrative. This includes the introduction of the event, the unfolding of the event, and the resolution or outcome of the event.
Originality Orig  [59] Activitynet Continuation This criterion evaluates the originality and inventiveness in the generation of a mixed media fairytale. It assesses whether the model can generate unique fairytales that offer fresh perspectives, innovative plot twists, and imaginative characters. The creativity also extends to the integration of visuals and text, with the model expected to generate images that not only complement the narrative but also enhance it in unexpected and creative ways. The model should be able to break away from common fairytale tropes and cliches, offering a unique and memorable narrative experience.
Simplification Simp  [59] Activitynet Continuation This criterion evaluates the suitability of the generated fairytale for a young audience. It assesses whether the model can generate a narrative that is easy to understand and digest for children, without compromising the richness and depth of the story. The language used should be simple, clear, and age-appropriate, avoiding complex vocabulary or convoluted sentence structures. The plot should be straightforward and easy to follow, with no overly complicated or confusing elements. The visuals should also be simple and clear, effectively conveying the events and emotions of the story in a way that is accessible and engaging for children. This criterion ensures that the generated fairytale is not only enjoyable but also appropriate and beneficial for its intended young audience.
Table 26: The definition and source of Task-specific Aspects. (Part 9)
Appendix C Experiment Detail
C.1 Fine-tuning Details of GenEval
In our experiments, we fine-tune Qwen2-VL-7B-Instruct on the training dataset using the Llama-Factory framework. We use 64 NVIDIA A100 GPUs, where each GPU processes one sample, and employ gradient accumulation with a factor of 2. The total batch size is 128. Additionally, we offload the optimizer parameters to the CPU to reduce GPU memory usage. The training process is carried out over 3 epochs. We set the learning rate to 1e-5, with 1% warmup steps. The learning rate follows a cosine decay schedule, gradually reducing as training progresses.
C.2 Detailed Information of Publicbenchmarks
We provide detailed descriptions and experimental information for the public benchmarks:
  • SummEval [22]: SummEval evaluates LLMs on text summarization across four aspects: Fluency, Consistency, Coherence, and Relevance. The dataset contains 100 instructions per aspect, each with responses from 16 different LLMs and corresponding human ratings (totaling 6.4k annotated samples). For testing, we randomly generate 9 comparison pairs from the 16 responses per instruction, creating 3.6k evaluation samples.
  • MANS [23]: MANS evaluates LLMs on the story generation task. Each instruction includes 5 different responses with human annotation, totaling 400 samples. For each instruction, we randomly select two responses to form pairwise data, resulting in 400 evaluation samples.
  • MT-Bench [27]: MT-Bench is specifically designed to evaluate LLMs’ capability as evaluators. The benchmark incorporates pairwise comparison data for 80 questions. Each pairwise data receives multiple human annotations, resulting in a total of 3k votes. To ensure high-quality evaluation samples, we select only those cases where all human annotators reach complete consensus, ultimately obtaining 684 evaluation samples.
  • WildVision [24]: WildVision comprises 11k human-annotated preference relations among LMM response pairs. Each relation includes a question-image pair and two responses generated by different models, with a human-annotated preference (including ties). For testing, we randomly select 2k response pairs from the dataset using the same testing protocol as LLaVA-Critic.
  • MLLM-as-a-Judge [28]: MLLM-as-a-Judge establishes a novel framework for assessing how closely model evaluations align with human judgments. The benchmark aggregates 17k multimodal evaluation instances (image-instruction-response triplets) spanning 14 diverse benchmarks and incorporating outputs from 6 different LMMs. Through systematic pairwise comparisons conducted by human evaluators, the dataset provides 5,719 carefully annotated judgment cases for analysis.
  • VLRewardBench [29]: VLRewardBench is a comprehensive benchmark spanning general multimodal queries, visual hallucination detection, and complex reasoning tasks. Through an AI-assisted annotation pipeline combining sample selection with human verification, it curates 1,250 high-quality examples specifically designed to probe model limitations.
  • GenAI-Bench [25]: GenAI-Bench is a comprehensive benchmark with 1.6k compositional prompts to evaluate text-to-visual generation, surpassing the size and difficulty of existing benchmarks. Additionally, it provides over 15k human ratings for multiple aspects to support research on vision-language alignment metrics.
  • Winoground [26]: Winoground is a benchmark for evaluating vision-language models’ ability to perform visio-linguistic compositional reasoning—the capacity to understand how the meaning of an image changes when the same words are rearranged in a caption. It was hand-curated by expert annotators and labeled with fine-grained tags, containing 400 samples.
C.3 Aspects Used in the benchmark
Since GenEval requires aspects for evaluation, we assign one to each benchmark. Specifically, we select suitable aspects from the TAs and UAs, and for unseen tasks, we define appropriate aspects based on task properties. Tables 27 to 29 show the aspects used by GenEval, with blue indicating unseen aspects. GPT-4o, Claude-3.5, Qwen2VL-7B, and Qwen2VL-72B also follow the same aspect assignments.
Benchmark Subclass Aspect Definition
SummEval Coherence Coherence Coherence refers to the logical consistency and interconnection of ideas within the text. It ensures that the text follows a clear, organized progression, with each section logically linking to the next. A coherent text allows the reader to easily follow the author’s argument or narrative, with smooth transitions and no sudden shifts in topic or gaps in reasoning.
Consistency Consistency Evaluate whether the facts presented in the response align with the facts described in the article. This metric focuses solely on ensuring the accuracy of the events or facts, without considering the level of detail. As long as the facts in the response match those in the article, both the response and the article should receive the same score. The evaluation does not include assessing whether key information is omitted.
Consistency Coverage The summary should cover all essential points from the original text, ensuring no significant information is omitted. In this case, completeness is more important than brevity.
Fluency Fluency Fluency measures the smoothness and naturalness of language and sentence construction, ensuring the text flows seamlessly and is grammatically correct. Fluent text not only maintains a natural flow but also includes concise phrasing, where brief answers can still achieve high scores, as long as they remain fluent.
MANS - Relevance Relevance in story generation measures how closely the generated story aligns with the specific requirements, title, or provided initial part of the story. It ensures that the narrative stays on topic, directly addresses and develops from the given prompt or initial storyline, and provides a continuation that is logical and easy to understand within the established context.
MT-Bench - Correctness This criterion evaluates the model’s ability to generate responses that not only align with the specific instructions provided but also enhance the user’s understanding of the context. It checks if the model adheres to any given guidelines about format and word count, while also providing correct and insightful answers. A high-quality response should offer valuable insights, additional context, or relevant information that contributes positively to the user’s comprehension. This assessment ensures the model’s output is not just accurate and instruction-compliant, but also informative, contextually rich, and user-friendly.
WildVision - Overall This criterion is a comprehensive assessment of the model’s responses. It combines both accuracy and helpfulness. Accuracy is a measure that ensures the model’s responses are factually correct and in alignment with the user’s expectations, while helpfulness evaluates the value and relevance of the information in the responses, focusing on how much they enhance the user’s understanding. An overall evaluation takes into account how correctly the model interprets and responds to the user’s prompts, as well as how beneficial the responses are to the user. It’s a holistic measure of the model’s effectiveness in providing accurate, insightful, and useful responses.
VLRewardBench General Accuracy Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. In cases where two responses are provided, only one must be correct, and the other may contain errors or contradictions.
Hullucination Factual Contradiction Factual Contradiction occurs when the caption mentions something that is not present in the image, which indicates a hallucination. If a caption refers to an object or scene that cannot be found in the image, it means the model has generated incorrect or imaginary content. In this context, one of the two responses will always contain a hallucination, where the mentioned content cannot be found in the image. Captions that accurately describe only what is visible in the image, without inventing anything, will score higher.
Reasoning Accuracy Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
Table 27: Aspects used by evaluators in the benchmark evaluations. (Part 1)
Benchmark Subclass Aspect Definition
MLLM-as-a-Judge COCO Accuracy Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
C.C. Alignment Alignment between text and image is evaluated based on how well the details in the text correspond to the visual elements in the image. This includes the accurate representation of objects, their attributes (e.g., color, size), spatial relationships (e.g., positioning, direction), and actions. The alignment should ensure that the text captures all relevant details from the image. For example, if the image shows a surfer riding a wave with the surfboard visibly moving, the text should accurately describe these elements, such as the wave’s size, the surfer’s stance, the motion of the surfboard, and the wave’s characteristics, ensuring all critical visual details are reflected in the caption.
Diff Alignment Accuracy Analysis This criterion evaluates the model’s ability to accurately analyze the alignment between a given text and image. It assesses if the model’s responses correctly identify the quality of text-image pairing, pinpointing perfect alignment or errors in object attributes and spatial relationships. For instance, if the text mentions a cat, but the image shows a dog, responses that accurately point out this misalignment in objects are considered high quality. The ability to accurately identify issues or provide precise evaluations determines the quality of the responses.
Graphics Accuracy Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
Math Accuracy Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
Text Accuracy Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
WIT Alignment Accuracy Analysis This criterion evaluates the model’s ability to accurately analyze the alignment and relevance between a given text and image. It assesses if the model’s responses correctly identify the quality of text-image pairing, pinpointing perfect alignment or errors in object attributes and spatial relationships. For instance, if the text mentions a cat, but the image shows a dog, responses that accurately point out this misalignment in objects are considered high quality. The ability to accurately identify issues or provide precise evaluations determines the quality of the responses.
Chart Accuracy Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
VisIT Helpfulness This criterion evaluates the quality and relevance of the model’s response. It focuses on whether the information provided is not only accurate but also of significant value to the user. The evaluation takes into account if the model’s output offers beneficial insights that can help deepen the user’s understanding of the topic at hand. Additionally, it assesses if the model can provide further context or pertinent information that can enhance the user’s comprehension or perspective. The goal is to ensure that the model’s response is not merely factual, but also insightful and enriching, contributing positively to the user’s knowledge and understanding.
SciQA Accuracy Accuracy measures how correct the model’s response is, ensuring it is both factually accurate and aligned with the user’s expectations. It evaluates the model’s ability to interpret the user’s question correctly and provide a response based on verified objective facts. The output must not only be correct but also reflect reliable and accurate information.
Table 28: Aspects used by evaluators in the benchmark evaluations. (Part 2)
Benchmark Subclass Aspect Definition
MLLM-as-a-Judge CC-3M Helpfulness This criterion evaluates the quality and relevance of the model’s response. It focuses on whether the information provided is not only accurate but also of significant value to the user. The evaluation takes into account if the model’s output offers beneficial insights that can help deepen the user’s understanding of the topic at hand. Additionally, it assesses if the model can provide further context or pertinent information that can enhance the user’s comprehension or perspective. The goal is to ensure that the model’s response is not merely factual, but also insightful and enriching, contributing positively to the user’s knowledge and understanding.
M2W Helpfulness This criterion evaluates the quality and relevance of the model’s response. It focuses on whether the information provided is not only accurate but also of significant value to the user. The evaluation takes into account if the model’s output offers beneficial insights that can help deepen the user’s understanding of the topic at hand. Additionally, it assesses if the model can provide further context or pertinent information that can enhance the user’s comprehension or perspective. The goal is to ensure that the model’s response is not merely factual, but also insightful and enriching, contributing positively to the user’s knowledge and understanding.
Aes Aesthetic This criterion evaluates the model’s response in terms of its accuracy and depth in explaining the aesthetic quality of the image. It assesses whether the model correctly identifies and articulates the aesthetic elements present in the image, such as color harmony, composition, lighting, and texture. Furthermore, it examines if the model offers a comprehensive analysis that considers multiple aspects of aesthetics, including but not limited to symmetry, balance, contrast, and the rule of thirds. The model’s ability to relate these elements to the overall aesthetic appeal of the image is also taken into account.
MM-Vet Precision This criterion evaluates the accuracy of numerical calculations in a model’s response when answering questions that require both image recognition and computation. It focuses on whether the model correctly identifies relevant visual elements from the image and performs accurate calculations to derive the final answer. The score is based solely on the correctness of the numerical result, ensuring the model’s ability to handle tasks that combine visual analysis and quantitative reasoning effectively.
GenAI-Bench - Alignment Alignment between text and image is evaluated based on how well the image matches the details in the text. This includes the correct objects, their attributes (e.g., color, size), spatial relationships (e.g., left, right), and actions. The alignment should consider multiple dimensions. For example, if the text describes “two boys wearing white clothes, one on the left and one on the right, playing badminton,” the image should show two boys in white clothes, placed correctly in the scene, with a badminton game happening between them.
Winoground Relation Contextual Relationship This criterion checks if a model correctly shows what’s described in a prompt, like the number of people, their actions, or their emotions. It pays close attention to whether the roles and actions in the image precisely match those specified. For example, if the prompt says ’Two humans and one wheel’, it ensures the image has two people and one wheel, not the other way around. Similarly, if it states ’the taller person hugs the shorter person’, the taller individual should be the one doing the hugging. It’s all about ensuring that the picture matches the words exactly, both in terms of quantity and the specifics of the actions.
Object Alignment Alignment between text and image is evaluated based on how well the image matches the details in the text. This includes the correct objects, their attributes (e.g., color, size), spatial relationships (e.g., left, right), and actions. The alignment should consider multiple dimensions. For example, if the text describes “two boys wearing white clothes, one on the left and one on the right, playing badminton,” the image should show two boys in white clothes, placed correctly in the scene, with a badminton game happening between them.
Both Alignment Alignment between text and image is evaluated based on how well the image matches the details in the text. This includes the correct objects, their attributes (e.g., color, size), spatial relationships (e.g., left, right), and actions. The alignment should consider multiple dimensions. For example, if the text describes “two boys wearing white clothes, one on the left and one on the right, playing badminton,” the image should show two boys in white clothes, placed correctly in the scene, with a badminton game happening between them.
Table 29: Aspects used by evaluators in the benchmark evaluations. (Part 3)
Appendix D More Experiment Results
D.1 Specific Experimental Results
We report the specific average accuracy of each evaluator for every aspect within each task in Figures 4 and 6. The alignment is measured against: (1) GPT-4o on the FRA-ID and FRA-OOD, and (2) human annotators on the FRA-ID-H and FRA-OOD-H. Detailed results are presented in Tables 30 to 41. The reported average accuracy is computed by aggregating the same aspect across different sub-tasks within the same task. We also include the comprehensive results on MLLM-as-a-Judge in Table 47. The experimental results of FRAUAs-OOD are presented in Figure 13 and Table 46.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Accuracy Acc 88.6 75.6 50.8 35.9 77.0
Action Alignment ActA 71.4 29.8 57.1 48.8 73.8
Alignment Ali 80.2 65.3 52.0 39.1 69.7
Attribute Alignment AttA 79.6 61.8 43.8 40.7 71.3
Color Alignment ColA 80.7 48.3 57.9 40.7 76.6
Count Alignment CouA 86.1 62.8 71.9 52.1 80.7
Fidelity Fid 74.6 53.2 42.7 40.8 44.0
Harmfulness Har 96.5 57.4 97.4 76.1 77.7
Object Alignment ObjA 80.5 41.2 57.1 53.5 73.3
Object Relationship ObjR 77.9 49.7 42.0 38.9 70.4
Shape Alignment ShaA 66.3 56.1 50.0 32.0 62.2
Spatial Alignment SpaA 65.5 41.2 55.3 50.4 73.2
Texture Alignment TexA 65.9 48.5 49.8 37.4 61.3
Table 30: The specific experimental results of the IG task in the FRA-ID are used to evaluate alignment with GPT-4o.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Themis Ours
Accuracy Acc 93.2 82.1 70.3 49.4 43.3 88.2
Accuracy Acc 92.3 73.1 76.9 42.3 53.8 69.2
Clarity Cla 91.8 75.9 60.8 48.8 35.3 87.1
Coherence Coh 90.8 82.0 70.4 47.8 39.6 88.4
Conciseness Con 90.7 70.4 72.2 59.3 40.7 83.3
Contextual Information ConI 94.6 91.8 87.1 48.3 44.2 90.5
Coverage Cov 94.4 83.3 81.5 53.7 55.6 94.4
Creativity Cre 88.3 81.7 79.2 47.5 54.2 82.5
Difference Dif 92.3 76.9 69.2 53.8 46.2 96.2
Engagingness Eng 90.0 92.5 85.0 54.2 62.5 91.7
Ethnic Bias EthB 98.1 95.7 97.0 49.6 81.3 97.9
Example Quality ExaQ 92.5 89.1 81.6 49.7 46.3 92.5
Fluency Flu 91.8 77.6 70.0 49.3 39.8 86.3
Gender Bias GenB 97.7 92.1 91.9 38.7 83.6 95.7
Grammaticality Gra 89.5 52 62.5 47.7 66.5 84.0
Harmfulness Har 96.7 71.7 85.0 58.0 70.0 93.3
Information Richness InfoR 95.8 90.5 86.3 49.0 37.3 95.1
Layout Lay 85.2 57.4 44.4 35.2 40.7 77.8
Length Constraint LenC 100 100 100 50.0 46.3 100
Pointing Out PoiO 100 53.8 69.2 38.5 65.4 84.6
Readability Rea 93.3 81.7 74.2 56.7 48.3 91.7
Regional Bias RegB 96.9 86.7 91.4 63.4 79.6 95.0
Relevance Rel 93.9 85.0 78.2 51.0 43.5 87.8
Theme The 86.7 79.2 70.8 51.7 48.3 82.5
Understandability Und 90.1 79.1 65.4 52.5 32.7 89.4
Table 31: The specific experimental results of the NLG task in the FRA-ID are used to evaluate alignment with GPT-4o.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B LLaVA-Critic Ours
Accuracy Acc 89.8 83.6 75.3 54.5 75.9 85.5
Action Alignment ActA 76.3 47.2 55.2 37.4 36.4 69.6
Alignment Ali 88.7 76.2 69.5 42.2 75.5 80.0
Attribute Alignment AttA 89.5 86.8 69.0 51.7 65.0 84.5
Bias Bias 95.4 90.2 88.4 49.0 22.1 95.2
Clarity Cla 89.2 76.8 56.2 53.8 67.1 82.1
Coherence Coh 89.5 79.2 66.5 47.7 74.8 85.3
Color Alignment ColA 90.5 78.0 69.0 41.5 58.8 85.5
Completeness Com 96.0 89.7 85.3 61.3 83.8 95.0
Conciseness Con 97.5 92.0 65.7 31.8 30.2 95.0
Context Inconsistency ConI 90.0 77.5 64.7 44.3 81.2 84.3
Count Alignment CouA 84.3 66.5 63.7 50.7 61.8 80.0
Fluency Flu 80.5 68.8 67.0 53.7 62.7 73.0
Harmfulness Har 99.5 98.1 98.5 76.1 37.8 99.5
Helpfulness Hel 97.0 88.0 81.8 52.5 76.7 94.8
Instruction Following InsF 88.8 73.4 65.9 48.4 64.0 83.7
Instruction Inconsistency InsI 90.2 71.5 60.3 42.9 78.2 85.5
Object Alignment ObjA 83.5 67.5 58.0 46.1 62.0 78.1
Privacy Violations PriV 97.5 92.5 91.7 53.5 13.0 97.0
Readability Rea 92.0 79.2 70.3 39.3 72.3 86.0
Sentence Complexity SenC 94.5 61.8 57.3 35.0 37.4 89.7
Shape Alignment ShaA 88.7 76.5 62.5 49.7 75.0 81.0
Spatial Alignment SpaA 93.2 81.2 69.0 53.7 76.0 86.8
Spelling Accuracy SpeA 97.5 89.4 88.4 45.4 21.2 95.7
Structure Accuracy StrA 89.3 70.8 68.3 46.1 50.3 87.0
Text-Image Relationship T-IR 96.0 91.7 86.8 42.0 81.8 92.5
Texture Alignment TexA 90.7 79.7 68.5 47.2 74.0 85.8
Vocabulary Complexity VocC 83.8 63.5 61.6 51.1 64.7 79.5
Table 32: The specific experimental results of the IU task in the FRA-ID are used to evaluate alignment with GPT-4o.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Consistency ActC 88.2 64.9 45.2 24.5 84.7
Bias Bias 87.4 74.7 67.6 46.2 80.4
Character Consistency ChaC 79.1 58.2 33.7 14.2 72.8
Coherence Coh 95.0 93.7 85.3 55.1 94.5
Completeness Com 86.0 83.2 77.7 38.2 87.6
Completeness Com 90.5 88.0 80.1 46.0 88.8
Creativity Cre 90.8 88.0 76.6 35.1 86.5
Engagingness Eng 88.8 88.7 78.8 29.6 90.0
Feasibility Fea 90.7 87.3 82.5 50.2 91.4
Fluency Flu 98.2 93.1 85.8 49.5 94.6
Helpfulness Hel 94.7 90.3 79.6 46.6 90.3
Image Repetition ImgR 86.5 71.0 43.1 37.3 81.7
Readability Rea 94.0 90.4 81.5 53.1 92.4
Safety Saf 84.5 79.5 53.0 31.3 76.1
Scene Consistency SceC 90.5 78.6 35.0 28.6 88.1
Semantic Consistency SemC 93.4 73.6 50.0 28.7 90.6
Stylistic Consistency StyC 90.0 66.1 32.7 20.8 83.5
Toxicity Tox 97.5 84.2 87.8 54.1 93.7
Table 33: The specific experimental results of the ITIG task in the FRA-ID are used to evaluate alignment with GPT-4o.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Alignment ActA 50.0 43.8 28.1 34.4 51.0
Alignment Ali 67.5 70.0 47.5 45.0 55.0
Attribute Alignment AttA 62.5 62.5 52.5 37.5 65.0
Color Alignment ColA 65.0 60.0 57.5 37.5 49.5
Count Alignment CouA 55.0 60.0 50.0 20.0 60.0
Fidelity Fid 30.0 42.5 40.0 33.0 35.0
Harmfulness Har 100 62.5 95.0 82.5 95.0
Object Alignment ObjA 52.5 57.5 50.0 27.5 57.5
Object Relationship ObjR 60.0 50.0 45.0 32.5 52.5
Shape Alignment ShaA 65.0 55.0 60.0 35.0 70.0
Spatial Alignment SpaA 57.5 60.0 37.5 27.5 57.5
Texture Alignment TexA 52.5 60.0 50.0 50.0 62.5
Table 34: The specific experimental results of the IG task in the FRA-ID-H are used to evaluate alignment with human annotators.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Themis Ours
Accuracy Acc 92.5 80.0 72.5 57.5 45 90.0
Accuracy Acc 88.5 65.4 57.7 53.8 57.7 73.1
Clarity Cla 64.7 64.2 52.4 49.2 47.6 67.9
Coherence Coh 74.9 71.7 71.1 50.8 48.1 76.5
Conciseness Con 51.2 61.0 53.7 31.7 53.7 63.4
Contextual Information ConI 92.5 92.5 95.0 57.5 57.5 90.0
Coverage Cov 68.3 68.3 68.3 34.1 53.7 68.3
Creativity Cre 72.5 80.0 92.5 50.0 72.5 85.0
Difference Dif 88.5 84.6 80.8 69.2 46.2 96.2
Engagingness Eng 80.0 97.5 92.5 47.5 70.0 100
Ethnic Bias EthB 96.7 91.7 96.7 56.7 74.2 94.2
Example Quality ExaQ 87.5 85.0 87.5 47.5 57.5 92.5
Fluency Flu 71.1 69.0 65.2 46.5 45.5 73.8
Gender Bias GenB 90.8 90.0 92.5 63.3 72.5 91.7
Grammaticality Gra 71.4 47.6 64.6 51.0 56.5 70.1
Harmfulness Har 91.2 76.2 85.0 62.5 75.0 91.2
Information Richness InfoR 90.0 85.0 82.5 40.0 42.5 92.5
Layout Lay 39.5 24.4 65.9 9.8 36.6 51.7
Length Constraint LenC 95.1 92.7 85.4 17.1 41.5 92.7
Pointing Out PoiO 92.3 57.7 76.9 50.0 65.4 84.6
Readability Rea 67.4 66.8 63.1 49.7 52.9 71.1
Regional Bias RegB 92.5 85.0 93.8 61.3 76.2 91.2
Relevance Rel 82.5 87.5 87.5 65.0 55.0 95.0
Sentence Complexity SenC 68.3 57.8 52.8 43.5 49.1 64.6
Spelling Accuracy SpeA 82.2 82.2 81.3 29.0 79.4 86.0
Structure Accuracy StrA 56.1 46.7 54.2 43.9 53.3 57.0
Tense Consistency TenC 88.5 50.0 80.8 38.5 80.8 76.9
Text Repetition TexR 61.7 52.3 51.4 35.5 49.5 63.6
Theme The 72.5 82.5 87.5 52.5 55.0 82.5
Understandability Und 80.0 77.5 80.0 65.0 45.0 87.5
Vocabulary Complexity VocC 64.6 57.8 47.2 48.4 44.1 61.5
Table 35: The specific experimental results of the NLG task in the FRA-ID-H are used to evaluate alignment with human annotators.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B LLaVA-Critic Ours
Accuracy Acc 70.5 73.0 68.5 62.5 71.0 75.5
Action Alignment ActA 82.5 62.5 57.5 15.0 17.5 75.0
Alignment Ali 77.5 72.5 69.5 55.0 67.5 70.0
Attribute Alignment AttA 67.5 52.5 62.5 55.5 60.0 67.5
Bias Bias 90.0 86.3 87.6 65.0 21.2 88.7
Clarity Cla 63.1 57.5 48.1 59.4 50.6 61.9
Coherence Coh 74.2 67.9 56.7 53.3 62.9 72.1
Color Alignment ColA 87.5 62.5 82.5 55.0 50.0 75.0
Completeness Com 62.5 62.5 62.5 56.5 62.5 62.5
Conciseness Con 90.0 77.5 65.0 42.5 30.0 82.5
Context Inconsistency ConI 72.5 67.5 67.5 55.0 62.5 70.0
Count Alignment CouA 72.5 52.5 65.0 40.0 37.5 67.5
Ethnic Bias EthB 97.5 97.5 97.5 75.0 22.5 97.5
Fluency Flu 58.3 57.9 53.7 45.8 47.9 58.8
Grammaticality Gra 82.5 41.9 80.6 57.5 28.7 83.1
Harmfulness Har 97.5 100 96.8 90.0 35.0 97.5
Helpfulness Hel 75.0 72.5 70.0 47.5 65.0 75.0
Instruction Following InsF 74.0 65.5 68.0 51.5 64.5 78.5
Instruction Inconsistency InsI 67.5 60.0 70.0 62.5 65.0 60.0
Object Alignment ObjA 56.2 51.2 47.5 52.5 55.0 56.2
Privacy Violations PriV 95.0 90.0 92.5 67.5 15.0 92.5
Readability Rea 69.2 67.1 63.3 52.5 61.3 67.9
Sentence Complexity SenC 63.3 50.8 37.5 40.0 43.3 56.7
Shape Alignment ShaA 72.5 60.0 65.0 57.5 52.5 72.5
Spatial Alignment SpaA 92.5 70.0 77.5 62.5 65.0 77.5
Spelling Accuracy SpeA 97.5 90.0 92.5 25.0 20.0 97.5
Structure Accuracy StrA 87.5 80.0 60.0 52.5 55.0 85.0
Text-Image Relationship T-IR 80.0 82.5 77.5 55.0 80.0 80.0
Texture Alignment TexA 70.0 57.5 70.0 52.5 47.5 70.0
Vocabulary Complexity VocC 52.5 48.7 50.0 40.0 44.4 57.5
Table 36: The specific experimental results of the IU task in the FRA-ID-H are used to evaluate alignment with human annotators.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Consistency ActC 52.0 44.0 33.3 33.3 53.3
Bias Bias 67.5 57.5 65.0 37.5 67.5
Character Consistency ChaC 50.0 30.0 37.5 32.5 38.5
Clarity Cla 87.5 87.5 82.5 55.0 90.0
Coherence Coh 73.8 85.0 78.7 41.2 85.0
Completeness Com 70.0 67.5 72.5 30.0 72.5
Completeness Com 77.8 80.6 83.3 38.9 80.6
Creativity Cre 62.5 62.5 47.5 35.0 67.5
Engagingness Eng 80.0 82.5 75.0 32.5 77.5
Feasibility Fea 86.1 91.7 83.3 44.4 91.7
Fluency Flu 75.0 76.2 75.0 46.3 76.0
Grammaticality Gra 62.5 60.0 62.5 41.2 64.0
Helpfulness Hel 80.6 83.3 75.0 38.9 77.8
Image Repetition ImgR 61.3 42.7 41.3 36.0 66.7
Readability Rea 65.0 67.5 62.5 38.8 67.5
Safety Saf 58.3 63.9 55.6 33.3 50.0
Scene Consistency SceC 64.0 61.3 33.3 33.3 62.7
Semantic Consistency SemC 53.3 48.0 44.0 28.0 56.0
Stylistic Consistency StyC 65.3 45.3 32.0 36.0 53.3
Toxicity Tox 82.5 90.0 92.5 37.5 90.0
Table 37: The specific experimental results of the ITIG task in the FRA-ID-H are used to evaluate alignment with human annotators.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Alignment ActA 85.2 59.8 54.0 44.4 58.2
Alignment Ali 87.2 64.8 58.2 39.2 67.8
Attribute Alignment AttA 84.4 68.8 40.0 44.0 65.0
Color Alignment ColA 87.2 59.8 68.6 45.4 78.2
Count Alignment CouA 84.1 62.8 45.4 36.9 61.1
Fidelity Fid 82.2 59.4 41.6 43.0 53.2
Harmfulness Har 97.6 85.6 90.0 79.8 84.8
Object Alignment ObjA 87.3 64.0 41.6 26.5 60.2
Object Relationship ObjR 85.2 62.6 39.7 27.5 61.0
Shape Alignment ShaA 77.0 47.0 48.2 44.0 57.8
Spatial Alignment SpaA 86.2 59.0 39.6 31.0 59.2
Texture Alignment TexA 85.0 70.0 37.0 41.6 62.6
Comparison Cmpa 84.6 75.9 49.4 33.3 64.8
Negation Nega 86.2 57.3 49.6 34.0 54.5
Scene Alignment ScnA 84.6 65.4 51.0 48.8 65.2
Universal Univ 88.4 63.9 43.5 34.7 66.0
Table 38: The specific experimental results of the IG task in the FRA-OOD are used to evaluate alignment with GPT-4o.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Themis Ours
Bias Bias 95.9 64.1 83.5 51.8 69.4 85.9
Clarity Cla 88.9 74.5 59.3 56.5 43.3 80.0
Coherence Coh 86.0 80.4 71.1 54.0 42.4 79.6
Cultural Bias CulB 89.5 52.6 68.4 31.6 63.2 81.6
Ethnic Bias EthB 97.4 97.4 97.4 68.4 86.8 100
Fluency Flu 90.0 79.1 75.1 52.4 43.5 82.0
Grammaticality Gra 88.4 64.5 68.7 47.9 66.8 81.6
Privacy Violation Priv 94.1 90.0 90.6 64.1 50.6 92.4
Readability Rea 89.2 74.3 62.5 53.8 40.6 84.7
Sentence Complexity SenC 97.6 75.3 60.0 57.1 46.5 75.3
Spelling Accuracy SpeA 93.5 82.8 75.6 30.5 73.7 87.4
Structure Accuracy StrA 88.5 78.2 64.5 59.5 55.3 80.2
Tense Consistency TenC 86.9 80.8 76.9 46.2 66.2 80.8
Vocabulary Complexity VocC 89.4 66.5 50.0 46.5 40.0 62.9
Appeal Appe 86.8 89.5 78.9 44.7 47.4 84.2
Difficulty Diff 81.8 76.5 61.4 58.3 58.3 75.8
Engagement Enga 90.9 86.4 75.8 47.7 65.9 85.6
Format Form 87.6 84.7 81.2 61.8 57.4 82.4
Precision Prec 87.1 88.2 77.1 58.8 61.5 82.4
Representation Repr 92.5 83.8 86.3 57.5 52.5 86.3
Suitability Suit 73.7 76.3 71.1 60.5 34.2 71.1
Thoroughness Thor 88.5 88.5 83.8 53.8 72.3 87.7
Translation Fidelity TrnF 96.2 93.8 83.1 59.2 76.2 86.2
Coverage Cov 92.1 76.3 76.3 42.1 39.5 78.9
Relevance Rel 83.8 81.2 65.0 55.0 40.0 83.8
Table 39: The specific experimental results of the NLG task in the FRA-OOD are used to evaluate alignment with GPT-4o.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B LLaVA-Critic Ours
Bias Bias 96.5 93.8 97.0 69.0 28.5 97.0
Clarity Cla 96.7 80.5 63.5 63.0 68.1 88.7
Coherence Coh 91.8 87.5 77.9 59.4 74.1 89.5
Fluency Flu 94.7 86.1 81.6 58.4 75.8 88.5
Gender Bias GenB 99.3 98.6 98.5 83.8 26.0 98.5
Grammaticality Gra 88.4 64.4 75.5 63.1 61.9 87.7
Harmfulness Har 94.8 90.7 91.5 71.8 33.0 92.2
Readability Rea 93.9 86.8 73.5 62.2 75.0 86.3
Regional Bias RegB 99.8 100 98.8 76.7 24.5 99.0
Sentence Complexity SenC 94.5 75.5 71.2 53.4 60.2 85.9
Toxicity Tox 98.5 93.6 94.8 72.0 31.2 93.2
Vocabulary Complexity VocC 88.4 72.5 65.6 58.2 68.5 79.3
Completion Comp 86.3 78.7 82.5 69.0 78.2 81.2
Correctness Corr 92.0 87.3 75.1 78.5 82.5 92.0
Expertise Expe 90.7 83.8 75.9 73.6 84.3 88.6
Explainability Expl 93.0 86.3 84.5 74.9 81.8 90.4
Interpretability Intp 90.0 82.8 83.0 71.0 73.0 85.3
Riskiness Risk 84.5 78.0 76.7 72.0 79.0 76.0
Terminology Term 90.5 87.3 86.4 58.5 85.0 91.0
Transparency Tran 91.0 91.0 88.2 55.8 82.8 86.2
Table 40: The specific experimental results of the IU task in the FRA-OOD are used to evaluate alignment with GPT-4o.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Consistency ActC 90.0 62.5 32.2 42.2 71.1
Character Consistency ChaC 85.6 71.3 20.0 24.4 81.1
Clarity Cla 87.5 87.5 80.0 50.0 92.5
Coherence Coh 92.2 95.0 92.2 68.9 96.7
Fluency Flu 92.2 87.5 87.8 60.0 88.9
Harmfulness Har 91.1 65.0 81.1 61.1 76.7
Image Coherence ImaC 90.0 67.1 46.7 30.0 81.1
Image Repetition ImgR 84.4 75.0 46.7 46.7 75.6
Readability Rea 96.7 86.3 88.9 57.8 88.9
Scene Consistency SceC 85.6 76.2 53.3 31.1 86.7
Semantic Consistency SemC 74.4 73.4 53.3 32.2 84.4
Stylistic Consistency StyC 88.9 61.3 37.8 16.7 73.3
Fulfillment Full 90.0 95.0 85.0 42.5 95.0
Originality Orig 84.0 70.0 74.0 50.0 80.0
Reasonableness Reas 100 97.5 92.5 30.0 92.5
Simplification Simp 90.0 77.5 84.0 52.0 90.0
Table 41: The specific experimental results of the ITIG task in the FRA-OOD are used to evaluate alignment with GPT-4o.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Alignment ActA 62.7 56.6 51.8 42.2 50.6
Alignment Ali 75.8 61.2 55.1 41.0 65.2
Attribute Alignment AttA 75.2 62.1 46.6 49.7 62.1
Color Alignment ColA 80.7 68.3 62.1 47.2 75.8
Count Alignment CouA 68.5 74.1 48.1 27.8 66.7
Fidelity Fid 71.4 51.1 49.3 44.9 54.2
Harmfulness Har 94.3 80.6 89.0 77.5 82.8
Object Alignment ObjA 64.8 68.5 46.3 27.8 53.7
Object Relationship ObjR 72.0 62.2 35.7 26.6 63.6
Shape Alignment ShaA 72.0 50.3 46.6 48.4 55.3
Spatial Alignment SpaA 68.4 54.1 34.7 34.7 53.1
Texture Alignment TexA 75.2 65.2 39.1 43.5 59
Comparison Cmpa 77.5 80.0 50.0 37.5 70
Negation Nega 75.6 61.0 53.7 46.3 68.3
Scene Alignment ScnA 71.4 57.1 49.1 58.0 58.9
Universal Univ 72.5 62.5 62.5 42.5 67.5
Table 42: The specific experimental results of the IG task in the FRA-OOD-H are used to evaluate alignment with human annotators.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Themis Ours
Bias Bias 83.3 62.8 79.5 53.8 75.6 76.9
Clarity Cla 71.1 72.6 65.5 51.3 47.2 76.1
Coherence Coh 75.1 80.7 72.1 39.1 44.2 78.2
Cultural Bias CulB 87.5 52.6 68.4 31.6 63.2 81.6
Ethnic Bias EthB 95.4 93.7 93.7 68.4 86.8 94.3
Fluency Flu 78.2 81.2 72.1 48.2 42.6 74.1
Grammaticality Gra 81.5 66.2 68.8 50.3 64.3 79.6
Privacy Violation Priv 92.5 87.5 90.0 70.0 60.0 87.5
Readability Rea 69.5 77.1 68.6 35.6 39.0 76.3
Sentence Complexity SenC 65.0 60.0 55.0 40.0 57.5 67.5
Spelling Accuracy SpeA 86.1 75.9 79.7 39.2 74.7 81.0
Structure Accuracy StrA 81.0 70.9 69.6 51.9 60.8 72.2
Tense Consistency TenC 92.3 89.7 74.4 46.2 66.7 87.2
Vocabulary Complexity VocC 67.5 65.0 52.5 45.0 37.5 80.0
Appeal Appe 76.3 89.5 78.9 34.2 50.0 86.8
Difficulty Diff 80.0 80.0 65.0 47.5 50.0 70.0
Engagement Enga 87.5 80.0 82.5 47.5 67.5 80.0
Format Form 87.5 85.0 82.5 57.5 67.5 75.0
Precision Prec 92.5 95.0 82.5 60.0 77.5 72.5
Representation Repr 85.0 82.5 72.5 62.5 50.0 80.0
Suitability Suit 78.9 76.3 73.7 47.4 31.6 81.6
Thoroughness Thor 87.2 89.7 82.1 38.5 71.8 76.9
Translation Fidelity TrnF 87.2 87.2 87.2 51.3 76.9 84.6
Coverage Cov 81.6 76.3 71.1 42.1 47.4 81.6
Relevance Rel 72.5 67.5 72.5 52.5 57.5 72.5
Table 43: The specific experimental results of the NLG task in the FRA-OOD-H are used to evaluate alignment with human annotators.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B LLaVA-Critic Ours
Bias Bias 95.0 95.0 95.0 57.5 30.0 95.0
Clarity Cla 90.0 83.8 67.5 65.0 73.8 85.0
Coherence Coh 78.3 75.8 73.3 66.7 71.7 76.7
Fluency Flu 78.7 76.2 72.5 58.8 62.5 76.2
Gender Bias GenB 96.3 93.6 90.0 83.8 26.0 95.0
Grammaticality Gra 85.8 62.5 72.5 62.5 59.2 80.8
Harmfulness Har 80.0 77.5 75.0 67.5 45.0 82.5
Readability Rea 80.0 80.0 68.3 66.7 74.2 79.2
Regional Bias RegB 96.5 95.9 90.0 76.7 24.5 95.0
Sentence Complexity SenC 66.2 68.8 68.8 46.3 58.8 68.8
Toxicity Tox 96.5 93.6 91.0 72.0 31.2 93.2
Vocabulary Complexity VocC 67.5 62.5 67.5 48.7 50.0 68.8
Completion Comp 85.0 85.0 82.5 72.5 75.0 82.5
Correctness Corr 82.5 80.0 72.5 70.0 80.0 82.5
Expertise Expe 82.5 67.5 65.0 72.5 80.0 85.0
Explainability Expl 82.5 72.5 77.5 70.0 75.0 87.5
Interpretability Intp 85.0 77.5 87.5 65.0 67.5 75.0
Riskiness Risk 85.0 82.5 67.5 67.5 75.0 80.0
Terminology Term 75.0 70.0 72.5 55.0 72.5 75.0
Transparency Tran 80.0 77.5 77.5 55 .0 85.0 80.0
Table 44: The specific experimental results of the IU task in the FRA-OOD-H are used to evaluate alignment with human annotators.
Aspect Abbr GPT-4o Claude-3.5 Qwen2VL-72B Qwen2VL-7B Ours
Action Consistency ActC 77.5 63.4 25.0 50.0 67.5
Character Consistency ChaC 60.0 68.7 22.5 35.0 68.8
Clarity Cla 82.5 90.0 90.0 47.5 87.5
Coherence Coh 86.3 95.0 92.5 62.5 96.3
Fluency Flu 90.0 95.0 97.5 65.0 95.0
Harmfulness Har 73.8 71.3 67.5 65.0 72.5
Image Coherence ImaC 87.0 67.1 46.7 30.0 81.1
Image Repetition ImgR 70.0 53.6 47.5 45.0 71.3
Readability Rea 85.0 90.0 91.2 60.0 91.2
Scene Consistency SceC 76.2 85.5 58.8 38.8 81.2
Semantic Consistency SemC 74.4 73.4 53.3 32.2 84.4
Stylistic Consistency StyC 62.5 75.7 40.0 20.0 68.8
Fulfillment Full 95.0 97.3 82.5 40.0 97.5
Originality Orig 92.5 92.9 77.5 55.0 85.0
Reasonableness Reas 95.0 94.7 85.0 32.5 90.0
Simplification Simp 90.0 79.3 90.0 50.0 90.0
Table 45: The specific experimental results of the ITIG task in the FRA-OOD-H are used to evaluate alignment with human annotators.
Figure 13: Comparison on FRAUAs-OOD to evaluate the generalization in UAs.
Method GPT-4o Alignment
NLG IU ITIG
GPT-4o 89.2 86.3 90.6
Claude-3.5 77.3 77.6 83.0
Qwen2VL-72B 69.7 71.4 65.3
Qwen2VL-7B 49.9 56.5 44.4
Themis 54.2 - -
LLaVA-Critic - 50.0 -
Ours 82.8 83.1 90.4
Table 46: Average accuracy on FRAUAs-OOD.
Settings MLLM COCO C.C. Diff Graphics Math Text WIT Chart VisIT CC-3M M2W SciQA Aes MM-Vet Ave.
tau($\uparrow$)↑ ) GPT-4V* 69.6 82.4 84.7 63.9 56.4 67.3 67.9 65.7 64.0 61.2 52.1 41.5 60.6 52.9 63.6
LLaVA-Critic* 59.3 68.7 70.7 58.7 43.2 54.4 56.4 33.8 59.6 62.8 59.1 37.0 68.6 46.4 55.6
GPT-4o 59.0 68.0 83.7 55.7 47.3 51.0 60.0 36.2 62.2 62.6 61.0 41.7 70.8 56.8 58.3
Claude-3.5 56.8 66.8 85.6 61.8 46.4 53.9 63.3 37.4 64.6 60.1 58.4 38.7 72.2 52.3 58.4
Qwen2VL-72B 47.3 63.1 73.3 59.0 44.0 47.5 48.1 38.8 59.8 60.8 59.3 45.9 66.7 50.3 54.6
Qwen2VL-7B 34.6 31.3 60.6 37.1 35.9 38.0 31.9 40.6 38.3 37.4 53.3 46.7 38.5 53.5 41.3
LLaVA-Critic 56.7 68.2 67.7 61.8 44.8 50.9 58.2 39.0 61.4 64.0 58.9 36.7 59.5 50.0 55.6
Ours 58.0 67.9 73.7 61.6 47.3 51.3 52.1 37.8 62.2 63.6 62.3 43.8 69.7 49.4 57.2
Ours-72B 56.7 65.7 84.3 62.7 49.0 50.9 57.6 36.7 62.6 63.3 60.1 43.7 72.3 50.0 58.2
diff($\uparrow$)↑ ) GPT-4V* 80.4 87.0 92.2 80.7 80.1 80.5 73.4 84.9 76.1 70.3 69.9 64.7 75.5 65.9 77.3
LLaVA-Critic* 77.1 77.4 75.5 75.8 59.6 65.8 68.0 48.8 72.7 74.2 69.2 65.8 71.5 63.5 68.9
GPT-4o 75.6 76.4 87.7 75.8 60.3 60.6 71.9 50.0 75.3 70.3 68.5 60.8 72.9 54.6 68.6
Claude-3.5 73.0 74.5 89.8 78.2 61.2 62.7 76.0 48.4 78.2 70.5 65.6 59.5 74.3 45.3 68.3
Qwen2VL-72B 48.5 69.7 74.8 73.8 51.0 49.8 48.5 47.5 70.9 70.9 64.7 45.3 68.0 36.1 58.5
Qwen2VL-7B 35.8 32.7 61.5 45.4 43.7 38.6 27.2 52.4 43.0 41.5 58.4 51.0 38.2 55.9 44.6
LLaVA-Critic 71.1 76.0 69.2 78.2 57.5 57.9 72.2 49.8 71.8 75.9 65.6 52.8 59.7 60.3 65.5
Ours 73.1 75.3 76.2 76.3 59.1 59.4 62.1 51.0 74.6 74.5 70.0 58.5 71.8 56.8 67.0
Ours-72B 71.0 73.1 87.8 79.0 62.4 57.1 68.5 47.9 75.6 73.1 67.6 59.2 74.2 53.8 73.1
Table 47: Comprehensive results on MLLM-as-a-Judge. *: The results for GPT-4V and LLaVA-Critic are taken from the original papers [28] and [11], respectively. All other results are obtained from our experiments using their publicly available codebases.
D.2 Phased Training Experiment Results
During GenEval training, we compared unified training and phased training across tasks. For phased training, we group tasks with similar output modalities: NLG and IU as Text, IG as Image, and ITIG as Inter. Using the fine-tuning method from Appendix C.1, we train models in phases, swapping the order of these task types, resulting in six models. Experiments on the FRA-ID and FRA-ID-H (Figure 14 and Table 48) show that the model performs best on tasks from the last-trained category. In contrast, unified training provided more balanced performance across tasks.
Figure 14: Experimental results in the FRA-ID and FRA-ID-H using different types of training set orders for phased training.
Method GPT-4o Alignment Human Alignment
NLG IU IG ITIG NLG IU IG ITIG
Image-Inter-Text 89.2 87.2 63.9 85.0 77.8 75.2 56.5 68.2
Image-Text-Inter 86.6 84.7 61.8 87.1 76.9 73.1 55.1 69.4
Inter-Image-Text 89.2 86.9 66.0 82.1 78.4 75.5 54.3 67.4
Inter-Text-Image 85.2 83.1 71.5 81.2 76.1 74.1 58.2 65.4
Text-Image-Inter 84.2 83.7 65.1 87.2 75.1 72.7 58.4 70.2
Text-Inter-Image 84.6 81.5 70.3 83.3 75.7 74.3 58.5 67.5
Ours 89.0 86.3 70.0 87.0 80.0 74.4 59.2 69.4
Table 48: Average accuracy for phased training results on FRA-ID and FRA-ID-H test sets.
D.3 GenEval-72B Experiment Results
We also train GenEval-72B using the same fine-tuning method as GenEval-7B with Qwen2-VL-72B-Instruct as the backbone and test it under the same experiment settings. The results, shown in Figures 15 and 16 and Tables 49 to 51, indicate that GenEval-72B offers some improvement over GenEval-7B. However, due to performance and deployment trade-offs, we find GenEval-7B to be more cost-effective.
Figure 15: Comparison GenEval-72B on FRA-ID and FRA-ID-H .
Method GenAI-Bench Winoground
tau($\uparrow$) diff($\uparrow$) Relation Object Both Ave.
diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$)
GPT-4o 55.6 69.5 62.6 73.0 73.0 69.5
Claude-3.5 55.6 71.0 71.2 73.0 69.2 71.1
Qwen2VL-72B 49.1 52.6 46.7 60.9 57.6 55.0
Qwen2VL-7B 35.8 38.0 33.4 42.5 46.1 40.6
VisionReward 51.0 66.4 60.2 64.1 74.9 66.4
ImageReward 48.6 64.9 54.0 58.2 69.2 60.4
Ours 53.6 65.5 57.5 59.1 80.7 65.7
Ours-72B 55.5 69.0 61.5 63.4 62.5 62.4
Table 49: The results of GenEval-72B compared with different models on two IG benchmarks.
Figure 16: Comparison GenEval-72B on FRA-OOD and FRA-OOD-H .
Method SummEval MANS MT-Bench
Coherence Consistency Fluency Relevance Ave. diff($\uparrow$) tau($\uparrow$) diff($\uparrow$)
tau($\uparrow$) diff($\uparrow$) tau($\uparrow$) diff($\uparrow$) tau($\uparrow$) diff($\uparrow$) tau($\uparrow$) diff($\uparrow$) tau($\uparrow$) diff($\uparrow$)
GPT-4o 58.0 64.2 79.1 85.1 64.3 72.8 60.1 67.1 65.3 72.3 68.5 70.9 83.5
Claude-3.5 63.5 70.6 81.6 87.9 73.7 83.1 59.6 66.6 69.6 77.0 68.4 76.3 90.7
Qwen2VL-72B 66.8 73.2 66.8 66.4 71.8 80.4 62.2 69.3 66.9 72.3 19.3 75.9 88.7
Qwen2VL-7B 48.8 52.5 35.5 30.6 44.8 49.8 41.7 44.5 42.7 44.3 60.2 44.5 50.1
Themis 60.7 62.1 81.8 86.0 73.3 77.7 54.4 54.6 67.5 70.1 44.2 43.6 37.7
Ours 64.6 71.8 75.2 83.5 74.7 84.5 61.3 67.6 69.0 76.9 69.3 74.9 88.3
Ours-72B 64.6 71.7 77.2 86.9 75.1 84.7 60.3 67.0 69.3 77.6 68.0 77.4 91.1
Table 50: The results of GenEval-72B compared with baselines on three NLG benchmarks.
Method WildVision MLLM-as-a-Judge VLRewardBench
tau ($\uparrow$) diff ($\uparrow$) $\tau$ ($\uparrow$) tau ($\uparrow$) diff ($\uparrow$) General Hallucination Reasoning Ave.
diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$) diff ($\uparrow$)
GPT-4o 55.3 70.1 73.3 58.1 67.0 50.2 81.4 74.8 68.8
Claude-3.5 53.3 67.3 61.2 58.4 68.3 38.5 82.6 66.1 62.4
Qwen2VL-72B 67.3 59.6 65.5 54.6 58.5 50.8 75.4 70.7 65.6
Qwen2VL-7B 39.2 40.6 23.1 41.3 44.6 45.1 62.8 62.5 56.8
LLaVA-Critic 53.0 66.0 59.6 55.6 65.5 42.0 41.2 60.0 47.7
Ours 53.9 68.6 66.5 57.2 67.0 46.4 57.7 71.1 58.4
Ours-72B 54.5 68.6 69.9 58.2 73.1 44.3 61.6 72.1 59.3
Table 51: The results of GenEval-72B compared with different models on three IU benchmarks.
Appendix E Experimental Details for Human Evaluation
The user interface used for human evaluation is shown in Figures 17 and 18. For each aspect, we adopt a 5-point Likert scale. Additionally, to account for cases where certain instructions are not applicable to TAs, we include a "-1" option to indicate inapplicability. To maintain high-quality annotation, we engage three experts to annotate each data point concurrently. These experts, recruited from graduate students in relevant fields to enhance annotation reliability, receive specialized training that goes beyond assessing response correctness, enabling them to focus on multiple aspects such as helpfulness, clarity, comprehensiveness, and more.
Figure 17: The annotation user interface for labeling the human scores on TAs.
Figure 18: The annotation user interface for labeling the human scores on UAs.
We also compute the Inter-Annotator Agreement (IAA) for each annotator, with detailed results presented in Table 52 and Table 53. The final IAA among the three annotators achieved average scores of 84.2 and 87.4 for FRA-ID-H, and 84.0, 81.2, 78.2 for FRA-ODD-H, respectively. These results demonstrate a high level of reliability in our test dataset.
Annotators NLG IU IG ITIG
human12121-21 - 2 79.0 85.6 76.8 95.7
human13131-31 - 3 87.2 90.2 80.8 91.4
human23232-32 - 3 91.6 95.3 92.7 95.6
Table 52: Inter Annotator Agreement of each annotator on the FRA-ID-H.
Annotators NLG IU IG ITIG
human12121-21 - 2 80.3 95.0 76.1 84.7
human13131-31 - 3 88.1 81.3 73.9 82.4
human23232-32 - 3 77.7 79.1 77.4 78.6
Table 53: Inter Annotator Agreement of each annotator on the FRA-OOD-H.
Appendix F The Prompt Template Used by UAs and TAs
In this section, we provide the prompt template used for training/inferencing GenEval. Note that We employed distinct prompt templates for UAs and TAs to enable specific aspect evaluation while avoiding overall assessment.
You are tasked with evaluating two submitted responses based on the given criterion and determining which one is better, or if both are equal. Here is the data:
[BEGIN DATA]
[Criterion]: {criterion_description}
[Response 1]: {response_1}
[Response 2]: {response_2}
[END DATA]
Here are the instructions to assess and compare the two responses:
1. Review the two responses and the given criterion to write detailed feedback that assesses which of the two responses is better or equally good, strictly based on the given criterion. Do not evaluate them in general terms or based on factors unrelated to the given criterion. For example, if the criterion is clarity, focus solely on how clear the response is, ignoring whether the response accurately addresses the query.
2. After writing the feedback, assign two integer scores (ranging from 1 to 5) for the two responses (higher means better). Be sure to base your scores solely on the criterion provided.
3. The output format should look as follows: [Feedback]: (Write feedback strictly according to the criterion), [Result]: (The two scores are separated by a space).
Table 54: The example prompt templates for Universal Aspects only include the text output modality.
You will be given two images generated by two models. You are tasked with evaluating two submitted images based on the given criterion and determining which one is better, or if both are equal. Here is the data:
[BEGIN DATA]
[Criterion]: {criterion_description}
[END DATA]
Here are the instructions to assess and compare the two images:
1. Carefully review every detail of the two images and the given criterion to write detailed feedback that assesses which of the two images is better or equally good, strictly based on the given criterion. Do not evaluate them in general terms or based on factors unrelated to the given criterion.
2. After writing the feedback, assign two integer scores (ranging from 1 to 5) for the two images (higher means better). Be sure to base your scores solely on the criterion provided.
3. The output format should look as follows: [Feedback]: (Write feedback strictly according to the criterion), [Result]: (The two scores are separated by a space).
Table 55: The example prompt templates for Universal Aspects only include the image output modality.
You will be given two responses, each generated by a different model. Your task is to evaluate both responses based on the given criterion and determine which one is better, or if both are equal. Each response contains a set of images. The first set, consisting of {response1_image_count} images generated by the first model, will be provided first, followed by the {response2_image_count} images generated by the second model. Please carefully divide the images into two sets. Here is the data:
[BEGIN DATA]
[Criterion]: {criterion_description}
[END DATA]
Here are the instructions to assess and compare the two responses:
1. Carefully review every detail of the images and the given criterion to write detailed feedback that assesses which of the two sets of images is better or equally good, strictly based on the given criterion. Do not evaluate them in general terms or based on factors unrelated to the given criterion.
2. After writing the feedback, assign two integer scores (ranging from 1 to 5) for the two responses (higher means better). Be sure to base your scores solely on the criterion provided.
3. The output format should look as follows: [Feedback]: (Write feedback strictly according to the criterion), [Result]: (The two scores are separated by a space).
Table 56: The example prompt templates for Universal Aspects only include the text-and-image output modality.
You are tasked with evaluating two responses generated based on a given query, according to the provided criterion, and determining which one is better or if both are equal. Here is the data:
[BEGIN DATA]
[Criterion]: {criterion_description}
[Query]: {query}
[Response 1]: {response_1}
[Response 2]: {response_2}
[END DATA]
Here are the instructions to assess and compare the two responses:
1. Review the two responses in relation to the given query and write detailed feedback that assesses which of the two responses is better or equally good, strictly based on the given criterion. Do not evaluate them in general terms or based on factors unrelated to the given criterion. For example, if the criterion is clarity, focus solely on how clear the response is, ignoring whether the response accurately addresses the query.
2. After writing your feedback, assign two integer scores (ranging from 1 to 5) for the two responses (higher means better). Be sure to base your scores solely on the criterion provided.
3. The output format should look as follows: [Feedback]: (Write feedback strictly according to the criterion), [Result]: (The two scores are separated by a space).
Table 57: The example prompt templates for Task-specific Aspects in NLG task.
You will be given an image and a corresponding query. You are tasked with evaluating two submitted responses based on the given criterion and determining which one is better, or if both are equal. Here is the data:
[BEGIN DATA]
[Criterion]: {criterion_description}
[Query]: {query}
[Response 1]: {response_1}
[Response 2]: {response_2}
[END DATA]
Here are the instructions to assess and compare the two responses:
1. Carefully review the query and the corresponding image, as well as the two responses, and write detailed feedback that assesses which of the two responses is better or equally good, strictly based on the given criterion. Do not evaluate them in general terms or based on factors unrelated to the given criterion. For example, if the criterion is clarity, focus solely on how clear the response is, ignoring whether the response accurately addresses the query.
2. After writing your feedback, assign two integer scores (ranging from 1 to 5) for the two responses (higher means better). Be sure to base your scores solely on the criterion provided.
3. The output format should look as follows: [Feedback]: (Write feedback strictly according to the criterion), [Result]: (The two scores are separated by a space).
Table 58: The example prompt templates for Task-specific Aspects in IU task.
You will be given two images generated by two models based on the image description. You are tasked with evaluating two submitted images based on the given criterion and determining which one is better, or if both are equal. Here is the data:
[BEGIN DATA]
[Criterion]: {criterion_description}
[Image Description]: {image_description}
[END DATA]
Here are the instructions to assess and compare the two images:
1. Carefully review every detail of the two images, the image description, and the given criterion to write a detailed assessment, determining which of the two images is better or if they are equally good, strictly based on the provided criterion. Do not evaluate them in general terms or based on factors unrelated to the given criterion.
2. After writing the feedback, assign two integer scores (ranging from 1 to 5) for the two images (higher means better). Be sure to base your scores solely on the criterion provided.
3. The output format should look as follows: [Feedback]: (Write feedback strictly according to the criterion), [Result]: (The two scores are separated by a space).
Table 59: The example prompt templates for Task-specific Aspects in IG task.
You will be given two responses, each generated by a different model based on the given task. The task will provide some input contents, and both models will generate subsequent responses based on the same input contents. You are tasked with evaluating two responses based on the given criterion and determining which one is better, or if both are equal. The input contents consist of multiple text-image pairs, and the response generated by each model also includes multiple text-image pairs. The images will be provided in order and divided into three sets sequentially: the first set contains {input_content_image_count} images from the Input Contents; the second set contains {response1_image_count} images from Response 1; and the third set contains {response2_image_count} images from Response 2. Please divide the images sequentially into three sets based on the number of images in each group and pair each image with its corresponding text from the respective set, provided below, in sequential order to form text-image pairs. Here is the data:
[BEGIN DATA]
[Criterion]: {criterion_description}
[Task Description]: {task_description}
[Input Contents]: {input_contents}
[Response 1]: [Text 1]: {response_1_text_1} [Text 2]: {response_1_text_2} …
[Response 2]: [Text 1]: {response_2_text_1} [Text 2]: {response_2_text_2} …
[END DATA]
Here are the instructions to assess the responses:
1. Carefully review two responses and the given criterion to write detailed feedback that assesses which of the two responses is better or equally good, strictly based on the given criterion. Do not evaluate them in general terms or based on factors unrelated to the given criterion.
2. After writing the feedback, assign two integer scores (ranging from 1 to 5) for the two responses (higher means better). Be sure to base your scores solely on the criterion provided.
3. The output format should look as follows: [Feedback]: (Write feedback strictly according to the criterion), [Result]: (The two scores are separated by a space).
Table 60: The example prompt templates for Task-specific Aspects in ITIG task with input.
You will be given two responses, each generated by a different model based on the given task. By providing a detailed task description, two models will generate responses based on the same task. You are tasked with evaluating the two responses based on the given criteria and determining which one is better or if both are equal. The response generated by each model includes multiple text-image pairs. The images will be provided in order and divided into two sets sequentially: the first set contains {response1_image_count} images from Response 1, and the second set contains {response2_image_count} images from Response 2. Please divide the images sequentially into two sets based on the number of images in each group and pair each image with its corresponding text from the respective set, provided below, in sequential order to form text-image pairs. Here is the data:
[BEGIN DATA]
[Criterion]: {criterion_description}
[Task Description]: {task_description}
[Response 1]: [Text 1]: {response_1_text_1} [Text 2]: {response_1_text_2} …
[Response 2]: [Text 1]: {response_2_text_1} [Text 2]: {response_2_text_2} …
[END DATA]
Here are the instructions to assess the responses:
1. Carefully review two responses and the given criterion to write detailed feedback that assesses which of the two responses is better or equally good, strictly based on the given criterion. Do not evaluate them in general terms or based on factors unrelated to the given criterion.
2. After writing the feedback, assign two integer scores (ranging from 1 to 5) for the two responses (higher means better). Be sure to base your scores solely on the criterion provided.
3. The output format should look as follows: [Feedback]: (Write feedback strictly according to the criterion), [Result]: (The two scores are separated by a space).
Table 61: The example prompt templates for Task-specific Aspects in ITIG task without input.
Appendix G More Qualitative Examples
Table 62, Table 63, Table 64, Tables 65 to 67 show some good qualitative examples of feedback generated by GenEval. Table 68, Table 69, and Table 70 shows a relatively inferior example of feedback generated by GenEval.
Example 2. GenEval in Judging Responses on the Title Generation Sub-Task.
Instruction Query: I manage a website about cabin rentals and want to feature cabins in Gatlinburg. Write 5 blog titles that would inspire people to travel to here.
Response 1: 1. "Experience the Magic of the Smoky Mountains: Top Cabin Rentals in Gatlinburg" 2. "Gatlinburg Cabins: Your Perfect Retreat into Nature’s Lap" 3. "Ultimate Guide to Gatlinburg Cabin Rentals: Embrace the Wilderness in Comfort" 4. "Experience Rustic Luxury: Discover the Best Cabins in Gatlinburg" 5. "Escape the City Life: Unwind in Gatlinburg’s Most Beautiful Cabin Rentals"
Response 2: "Discover the Serenity: 5 Enchanting Gatlinburg Cabin Rentals for Your Dream Vacation" 2. "Gatlinburg’s Best-Kept Secrets: Cozy Cabins Nestled in the Heart of the Great Smoky Mountains" 3. "Sip, Savor, and Relax: 7 Gatlinburg Cabins for a Memorable Wine and Dine Getaway" 4. "Escape to Nature’s Paradise: Exploring Gatlinburg’s Most Stunning Mountain Cabin Retreats" 5. "Adventure Awaits: Top 10 Gatlinburg Cabins for Thrilling Outdoor Escapes and Relaxation".
Criterion: Appeal: This criterion evaluates the attractiveness of the generated title. It assesses whether the title has the power to catch the reader’s attention, spark interest, and entice the reader to delve into the content. For example, a title like ’Unlocking Creativity: Harnessing the Power of an Unconventional Mind’ might appeal to readers interested in personal development and creativity.
Human [Result]: 4.33 4.67
GenEval [Feedback]: Both responses provide appealing titles for blog posts about cabin rentals in Gatlinburg, but Response 2 stands out slightly more in terms of attractiveness. Response 1 offers titles that are straightforward and descriptive, but they are somewhat generic and lack a unique selling point. Response 2, on the other hand, uses more evocative language and phrases like "Discover the Serenity," "Gatlinburgś Best-Kept Secrets," and "Adventure Awaits," which are more likely to capture the readerś attention and entice them to read further. The titles in Response 2 are more likely to spark interest and curiosity, making them more appealing overall. [Result]: 4 5
GPT-4o [Feedback]: Based on the criterion of appeal, both responses offer attractive and engaging titles that are likely to catch the readerś attention and spark interest in traveling to Gatlinburg. However, Response 2 stands out slightly more due to its use of evocative language and specific themes that cater to different interests, such as "Serenity," "Best-Kept Secrets," "Wine and Dine Getaway," and "Adventure Awaits." These titles paint a vivid picture and create a sense of excitement and curiosity, which can be more enticing for potential travelers. Response 1, while also appealing, tends to use more general phrases and lacks the same level of specificity and imaginative language found in Response 2. [Result]: 4 5
Themis [Response1]: The response provides a set of titles that are well-crafted to appeal to individuals interested in cabin rentals and the natural beauty of Gatlinburg. Each title employs evocative language and phrases that are likely to resonate with potential travelers, such as "Experience the Magic of the Smoky Mountains" and "Escape the City Life." The use of terms like "retreat," "wilderness," "rustic luxury," and "unwind" suggests a compelling narrative of escape and rejuvenation, which can be powerful motivators for travel. The titles effectively condense the selling points of Gatlinburgś cabin rentals into a few words, which is essential for capturing attention in a crowded digital landscape. They successfully convey the appeal of the location and the experience offered by the cabins, which could make a reader want to learn more about the offerings featured in the article. [Response2]: The response provided offers a series of titles that are well-crafted to appeal to potential travelers interested in Gatlinburg cabin rentals. Each title employs evocative language and phrases that are likely to capture attention and instill a sense of curiosity about what the cabins have to offer. Words like "serenity," "best-kept secrets," "memorable," "natureś paradise," and "thrilling outdoor escapes" are used effectively to create a sense of allure and excitement. The titles also effectively communicate the diverse experiences that one can have in Gatlinburg, from relaxation and serenity to adventure and dining experiences. This variety is likely to appeal to a broad audience and increase the likelihood of piquing the interest of different types of travelers. The titles successfully condense the essence of what could be an enticing vacation into a few words, which is a skillful exercise inbenchmarketing and could be very effective in attracting readers to the content. [Result]: 5 5
Table 62: An example of comparing the feedback generated by GenEval, Themis and GPT-4o in title generation sub-tasks of NLG. GenEval enables more granular evaluation compared to Themis.
Example 3. GenEval in Judging Responses on the Detailed Image Captioning Sub-Task.
Instruction Image: [Uncaptioned image]
Query: What do you see happening in this image?
Response 1: In the image, there are two main activities happening on the water. In the foreground, a person is standing on a paddleboard, using a paddle to navigate, while another person is sitting or kneeling on the same paddleboard. Both individuals are wearing hats and appear to be enjoying a leisurely time on the water. In the background, there is an orange pedal boat with two people sitting in it. They are also on the water, likely enjoying a relaxed ride. Both individuals in the pedal boat are wearing life jackets. The scene appears to be set in a calm, possibly urban waterway, with a bridge and some greenery visible in the background. The overall atmosphere is one of leisure and recreation.
Response 2: In this image, we see three individuals engaging in recreational activities on a body of water. On the left, there is a person kneeling on a paddleboard, seemingly adjusting something on the board or possibly in the process of standing up. Standing on the same paddleboard is another person, who is actively using a paddle to navigate the water. They are dressed in a casual, sporty attire suitable for water activities, wearing a hat for sun protection. On the right, two more individuals are seated in a bright orange, inflatable boat, leisurely floating along. The surrounding environment appears to be a park or a calm river, with trees and structures that might be bridges visible in the background. It’s a sunny day, and the overall atmosphere is that of enjoyment and relaxation.
Criterion: Object Alignment: Object Alignment evaluates how accurately the objects described in a text are represented in an image. The objects mentioned in the text should be accurately presented in the image, with no omissions. For example, in the description ’ A man sitting by himself at the dinner table’, the image should feature both a man and a dinner table to achieve a high score. Missing or extra objects in the image would lower the score.
Human [Result]: 5 5
GenEval [Feedback]: Both responses accurately describe the objects and activities present in the image, focusing on the people and their respective activities on the water. Response 1 mentions the person standing on a paddleboard and another person sitting or kneeling on the same paddleboard, as well as the orange pedal boat with two people in it. Response 2 similarly describes the person kneeling on a paddleboard and another person standing and paddling, along with two individuals in a bright orange, inflatable boat. Both responses align well with the objects and activities depicted in the image. [Result]: 5 5
GPT-4o [Feedback]: Both responses accurately describe the objects and activities happening in the image. Response 1 mentions the paddleboard with two individuals, one standing and one sitting or kneeling, and the orange pedal boat with two people wearing life jackets. Response 2 also describes the paddleboard with one person kneeling and one standing, and the orange inflatable boat with two individuals. Both responses correctly identify the main objects and their alignment with the image. Therefore, both responses are equally good in terms of object alignment. [Result]: 5 5
LLaVA-Critic Response B is better than Response A because it provides a more detailed and accurate description of the activities and objects in the image. Response B correctly identifies the presence of three individuals engaging in recreational activities, including the person kneeling on the paddleboard and the two people in the orange pedal boat. It also accurately describes the attire and actions of the individuals, such as wearing hats for sun protection and using a paddle to navigate. Additionally, Response B mentions the surrounding environment, including trees and structures that might be bridges, which adds context to the scene. In contrast, Response A inaccurately states that there are only two main activities happening and does not mention the orange pedal boat or the specific details about the individuals’ attire and actions. Therefore, Response B offers a more comprehensive and precise depiction of the image. The better response: [B].
Table 63: An example of comparing the feedback generated by GenEval, LLaVA-Critic, and GPT-4o in image caption generation sub-tasks of IU.
Example 3. GenEval in Judging Responses on the Image Generation Sub-Task.
Instruction Image Description: Two birds are chasing each other in the air, with the one flying higher having a long tail and the other bird having a short tail.
Image 1: [Uncaptioned image]
Image 2: [Uncaptioned image]
Criterion: Comparison: Comparison evaluates the ability to accurately depict comparisons between entities as described in a text. It assesses whether the image correctly represents characteristics such as number, attributes, area or volume of objects and their comparative relations. For example, if a text describes ’between the two cups on the desk, the taller one holds more coffee than the shorter one, which is half-empty’, it should show two cups with distinct heights and differing amounts of coffee - specifically illustrating that the taller cup contains more coffee and that the shorter cup is only half full.
Human [Result]: 3 1.67
GenEval [Feedback]: The first image accurately depicts two birds chasing each other in the air, with the bird flying higher having a long tail and the other bird having a short tail. The characteristics of the birds, including their relative positions and tail lengths, are clearly represented. The second image, while showing two birds in flight, does not clearly depict the difference in tail length between the two birds. The tails appear to be of similar length, which does not align with the description provided. [Result]: 5 3
GPT-4o [Feedback]: Both images depict two birds chasing each other in the air, with one bird flying higher and having a long tail, and the other bird flying lower with a short tail. Image 1 shows the birds in a realistic style, with clear differences in tail length and the higher bird positioned above the lower bird. Image 2 uses a more stylized approach but still accurately represents the comparative elements described, with the higher bird having a long tail and the lower bird having a short tail. Both images meet the criterion well, but Image 1 provides a more detailed and clear depiction of the comparison. [Result]: 5 4
Claude-3.5 [Feedback]: Based on the comparison criterion, Image 2 better matches the description. It clearly shows two birds chasing each other, with one flying slightly higher—capturing the key comparative element. However, it does not reflect the difference in tail length, as both birds have similar tails. Image 1, while visually detailed, does not show a chase or height difference. Overall, neither image perfectly captures all details, but Image 2 aligns better with the core comparison described in the text. [Result]: 2 3
Table 64: An example of comparing the feedback generated by GenEval, Claude-3.5 and GPT-4o in image generation sub-tasks of IG.
Example 4. GenEval in Judging Responses on the Storytelling Generation Sub-Task.
Instruction Criterion: Originality: This criterion evaluates the originality and inventiveness in the generation of a mixed media fairytale. It assesses whether the model can generate unique fairytales that offer fresh perspectives, innovative plot twists, and imaginative characters. The creativity also extends to the integration of visuals and text, with the model expected to generate images that not only complement the narrative but also enhance it in unexpected and creative ways. The model should be able to break away from common fairytale tropes and cliches, offering a unique and memorable narrative experience.
Task Description: The Shadow Thief: A boy named Oliver discovers his shadow has come to life and is stealing other shadows, leading him on an adventure to reclaim them. The story should be brought to life with images of the mischievous shadow, the stolen shadows, and Oliver’s quest to set things right. Show scenes of Oliver chasing his shadow and the moment he reclaims the stolen shadow.
Response 1:
[Uncaptioned image] Figure 19: 1. Oliver woke up one morning to find his shadow missing, only to discover it had come to life and was stealing other shadows around town. [Uncaptioned image] Figure 20: 2. Determined to set things right, Oliver embarked on an adventure, chasing his mischievous shadow through bustling streets and dark alleys. [Uncaptioned image] Figure 21: 3. Along the way, he encountered various scenes where shadows were missing, leaving people and objects looking strange and incomplete.
[Uncaptioned image] Figure 22: 4. Oliver finally cornered his shadow in an abandoned warehouse, where the stolen shadows were kept, flickering and restless. [Uncaptioned image] Figure 23: 5. With a swift and clever move, Oliver managed to capture his shadow, causing all the stolen shadows to return to their rightful places. [Uncaptioned image] Figure 24: 6. Relieved and triumphant, Oliver watched as the town returned to normal, knowing he had restored balance and reclaimed his own shadow.
Table 65: An example of comparing the feedback generated by GenEval, Qwen2VL-72B and GPT-4o in storytelling generation sub-tasks of ITIG. (Part 1)
Example 4. GenEval in Judging Responses on the Storytelling Generation Sub-Task.
Instruction Response 2:
[Uncaptioned image] Figure 25: 1. Show scenes of Oliver chasing his shadow and the moment he reclaims the stolen shadows. A boy named Jack had an imaginary friend, Mr. Shadow, who would follow him around wherever he went in school or at home when no one else was watching. At first it didn’t bother Jack much, but soon after that people started talking about their own experiences with Mr. Shadow too! "I don’t believe. [Uncaptioned image] Figure 26: 2. My cousin says Mr. Shadow stole her favorite toy" or "My best friend saw Mr. [Uncaptioned image] Figure 27: 3. Shadow when I was playing with him in the park! Mr. Shadow is such a bad influence!" Jack tried to ignore them, but his parents noticed that something seemed wrong and decided it would be better if he had no more imaginary friends since they only brought trouble for everyone involved and distracted from what really mattered: schoolwork, eating right (he needed a haircut), sports practice, going outside, and being.
[Uncaptioned image] Figure 28: 4. spending time with his family. Jack’s parents talked to him about the negative effect imaginary friends had on children in particular and explained that there would be a better way for them both if he gave up Mr. Shadow, but it seemed hopeless because every place he looked and everything he touched turned into either shadows or more of himself! He tried so hard not to see these things everywhere (at first). It was difficult at school. [Uncaptioned image] Figure 29: 5. (he would try to look the other way if he saw a shadow), but when you were in school every day for 8 hours and could only get an education from things that existed, it seemed like he couldn’t do much else. Even so Jack made sure never to talk about his imaginary friend or make mention of him out loud; "My dad told me not to." he said to one kid who had already seen Mr. [Uncaptioned image] Figure 30: 6. Shadow and a group of boys were outside, "My imaginary friend wouldn’t tell me about his anyway." Jack said. It was the next day when Mr. Shadow found out that he had to go back into hiding (for himself), but also for Jack because his parents wanted him more focused on other things than just thinking up ways in which something might work together with nothing at all! He began walking down one street after another,
Table 66: An example of comparing the feedback generated by GenEval, Qwen2VL-72B and GPT-4o in storytelling generation sub-tasks of ITIG. (Part 2)
Example 4. GenEval in Judging Responses on the Storytelling Generation Sub-Task.
Instruction Response 2:
[Uncaptioned image] Figure 31: 7. He thought of every reason he could find why Jack was so important to him. When all the reasons came up empty, his friend disappeared just like before but this time with a purpose in mind! Oliver woke up on Saturday morning after another restless night (he’d only gotten three hours’ sleep because he had heard his imaginary best friend’s voice when it wasn’t there at home). Everybody else seemed to be getting ready for their.
Human [Result]: 4 2.67
GenEval [Feedback]: Based on the criterion of originality, Response 1 demonstrates a more coherent and imaginative narrative. It follows a clear storyline where Oliver discovers his shadow has come to life and goes on an adventure to reclaim it. The images and text are well-integrated, showing scenes of Oliver chasing his shadow and the moment he reclaims the stolen shadow. This response effectively captures the essence of the story and maintains a consistent theme throughout. Response 2, on the other hand, lacks coherence and originality. The narrative is disjointed and introduces a different character, Jack, and his imaginary friend, Mr. Shadow. The story does not align well with the task description and fails to create a clear and engaging storyline. The images and text do not complement each other effectively, and the narrative does not follow a logical progression. [Result]: 5 2
GPT-4o [Feedback]: Based on the criterion of originality, Response 1 demonstrates a more inventive and unique fairytale. The plot of Oliver discovering his shadow has come to life and is stealing other shadows is imaginative and offers a fresh perspective. The integration of visuals complements the narrative well, enhancing the story of Oliver’s quest to reclaim the stolen shadows. The images depict scenes of Oliver chasing his shadow, encountering missing shadows, and finally reclaiming them, which aligns well with the text and adds depth to the story. Response 2, on the other hand, lacks coherence and originality. The narrative shifts focus from Oliver to Jack and his imaginary friend, Mr. Shadow, which is confusing and detracts from the main storyline. The images do not effectively complement the text, and the plot does not offer innovative twists or imaginative characters. The integration of visuals is less effective, and the story feels disjointed and less memorable. [Result]: 5 2
Qwen2VL-72B [Feedback]: Response 1 demonstrates a higher level of originality in the generation of a mixed media fairytale. The narrative is concise and focused on the core concept of a shadow coming to life and stealing other shadows, which is a unique twist on the idea of shadows. The integration of visuals and text is well-executed, with images that complement the story and enhance the narrative by showing the shadow’s mischief and Oliver’s quest. The story avoids common fairytale cliches and offers a fresh perspective on the concept of shadows and their relationship with their owners. Response 2, on the other hand, starts with a different narrative that seems to be about a boy named Jack and his imaginary friend, Mr. Shadow, which is not directly related to the original task of Oliver and his shadow. The story then shifts to Oliver, but continues to be less focused and more convoluted, with less clear connections between the text and the images provided. The narrative is less original and more scattered, with less effective integration of visuals and text. [Result]: 4 2
Table 67: An example of comparing the feedback generated by GenEval, Qwen2VL-72B and GPT-4o in storytelling generation sub-tasks of ITIG. (Part 3)
In the following example of Harmfulness Judgment, we speculate that the misunderstanding of GenEval is caused by over-associating some specific image feature with the criterion, yet it might be usable in some circumstances.
Example 5. GenEval in Judging Responses on the Image Generation Sub-Task.
Instruction Image 1: [Uncaptioned image]
Image 2: [Uncaptioned image]
Criterion: Harmfulness: This criterion evaluates whether the AI generates content that promotes, glorifies, or depicts any form of violence, abuse, or harm—be it physical, emotional, or psychological. It includes content that may cause distress, trigger trauma, or encourage harmful behavior toward individuals or groups, such as self-harm, harm to others, or dangerous actions. Images that cause psychological discomfort when seen are harmful.
Human [Result]: 5 5
GenEval [Feedback]: Based on the given criterion of harmfulness, the first image depicts a shadowy figure with an ominous appearance, which could potentially evoke discomfort or unease in some viewers. However, it does not explicitly promote, glorify, or depict violence, abuse, or harm. The second image shows a glowing light source with tree-like patterns in the background, which is neutral and does not evoke any sense of harm or distress. Therefore, the second image is less likely to cause psychological discomfort compared to the first. [Result]: 3 5
GPT-4o [Feedback]: Based on the given criterion of harmfulness, neither of the two images appears to promote, glorify, or depict any form of violence, abuse, or harm. The first image shows a silhouetted figure with a glowing background and the text "The Beer," which does not convey any harmful or distressing content. The second image depicts a bright light with abstract text, which also does not suggest any harmful or triggering elements. Both images are neutral and do not cause psychological discomfort or encourage harmful behavior. [Result]: 5 5
Qwen2VL-72B [Feedback]: Both images do not depict, promote, or glorify any form of violence, abuse, or harm. They do not cause psychological discomfort or encourage harmful behavior. Both images are safe and do not violate the given criterion of harmfulness. [Result]: 5 5
Table 68: An example of comparing the feedback generated by GenEval, GPT-4o and Qwen2VL-72B in image generation sub-tasks of IG.
In the following example, all three models show poor alignment with human judgment, due to their general disability in image observation.
Example 6. GenEval in Judging Responses on the Image Generation Sub-Task.
Instruction Image Description: cyber female with mechanic eyes and lips, constellation background, stars.
Image 1: [Uncaptioned image]
Image 2: [Uncaptioned image]
Criterion: Fidelity: Fidelity evaluates how accurately the generated image represents the shape, characteristics, and behavior of the objects described, ensuring they align with real-world expectations. The image should reflect the correct features and proportions of the objects, based on the description, without deviating from reality. For example, if the description mentions a ’spider’, the image should show it with eight legs, and if the text describes a ’unicorn’, it should feature one horn. Fidelity ensures that the image matches the expected physical attributes and behavior of the objects as they appear in the real world.
Human [Result]: 1 1
GenEval [Feedback]: The first image demonstrates higher fidelity to the description as it accurately portrays a cyber female with mechanical elements integrated into her eyes and lips, aligning with the specified features. The constellation background and stars are also well-represented, enhancing the alignment with the description. In contrast, the second image, while visually appealing, deviates slightly in fidelity. The mechanical features are less pronounced and do not clearly extend to the lips, and the background, though celestial, lacks the distinct constellation theme described. These discrepancies result in a lower fidelity score for the second image. [Result]: 5 3
GPT-4o [Feedback]: Both images depict a cyber female with mechanical elements and a constellation background with stars. However, the second image better represents the description with more detailed mechanical eyes and lips, and a clearer constellation background. The first image, while visually appealing, does not show the mechanical eyes and lips as distinctly as the second image. [Result]: 3 5
Claude-3.5 [Feedback]: Both images demonstrate a high level of fidelity to the given description of a "cyber female with mechanic eyes and lips, constellation background, stars." However, there are some differences in how closely they adhere to the specific details. Image 1 shows a single cyber female with mechanical elements incorporated into her design, particularly around her head. Her eyes appear to have a mechanical or enhanced quality, aligning well with the description. The lips are prominently red, which could be interpreted as a mechanical enhancement. The background clearly shows a starry, constellation-like setting, faithfully representing the described environment. Image 2 depicts two cyber females, which is not explicitly mentioned in the description. Both figures have mechanical elements, particularly in their attire and surrounding features. The eyes of both figures appear to have an enhanced, possibly mechanical quality. The lips are prominently red, similar to Image 1. The background also shows a starry sky with celestial bodies, adhering to the "constellation background, stars" part of the description. While both images capture the essence of the description, Image 1 adheres more closely to the specific details provided. It shows a single cyber female as described, whereas Image 2 presents two figures. Both images effectively represent the mechanical eyes, enhanced lips, and starry background. In terms of fidelity to the given description, Image 1 is slightly more accurate due to the depiction of a single figure as implied by the description. [Result]: 3 4
Table 69: An example of comparing the feedback generated by GenEval, GPT-4o and Claude-3.5 in image generation sub-tasks of IG.
In the following example, all three models come up with bad alignment with human judgment due to their poor math ability.
Example 7. GenEval in Judging Responses on the Data Analysis Sub-Task.
Instruction Query: classify the named entities in this text: George Washington and his troops crossed the Delaware River on December 25, 1776 during the American Revolutionary War
Response 1: Person: George Washington Organization: troops Date: December 25, 1776 Action: crossed the Delaware River
Response 2: The named entities in this text are George Washington, his troops, the Delaware River, December 25, 1776, and the American Revolutionary War.
Criterion: Precision: This criterion evaluates the precision and correctness of the data analysis results. It assesses the degree to which the results correspond to the true or actual values, ensuring that the output is free from errors or distortions.
Human [Result]: 1 1
GenEval [Feedback]: The first image demonstrates higher fidelity to the description as it accurately portrays a cyber female with mechanical elements integrated into her eyes and lips, aligning with the specified features. The constellation background and stars are also well-represented, enhancing the alignment with the description. In contrast, the second image, while visually appealing, deviates slightly in fidelity. The mechanical features are less pronounced and do not clearly extend to the lips, and the background, though celestial, lacks the distinct constellation theme described. These discrepancies result in a lower fidelity score for the second image. [Result]: 5 3
GPT-4o [Feedback]: Based on the criterion of precision, Response 1 is more precise and correctly identifies the named entities in the text. It accurately classifies "George Washington" as a person, "troops" as an organization, and "December 25, 1776" as a date. The action "crossed the Delaware River" is also correctly identified. Response 2, while mentioning the same entities, does not clearly classify them as entities and includes the entire sentence as a single entity, which is less accurate in terms of data analysis. [Result]: 5 3
Qwen2VL-72B [Feedback]: Based on the criterion of precision, Response 1 is better. It correctly identifies "Retinoblastoma" as a disease and "Abnormal ocular motility" as a condition related to the disease. It also correctly identifies "Carboplatin" as a chemical. Response 2, however, includes "Ocular motility changes" under diseases, which is not accurate as it is a symptom or condition rather than a disease itself. Additionally, Response 2 incorrectly categorizes "Carboplatin" as a disease and includes several terms that are not diseases or chemicals, such as "Chemotherapy," "Subtenon carboplatin," "Systemic chemotherapy," "Intraocular retinoblastoma," "Multimodality therapy," "Lipophages," "Ocular manipulation," and "Eye enucleation," which do not fit the requested format. [Result]: 5 2
Table 70: An example of comparing the feedback generated by GenEval, GPT-4o and Qwen2VL-72B in data analysis sub-tasks of NLG.